### v1

# Cross-Entropy Games for Language Models: From Implicit Knowledge to General Capability Measures

Abstract.
Large Language Models (LLMs) define probability measures on text. By considering the implicit knowledge question of what it means for an LLM to know such a measure and what it entails algorithmically, we are naturally led to formulate a series of tasks that go beyond generative sampling, involving forms of summarization, counterfactual thinking, anomaly detection, originality search, reverse prompting, debating, creative solving, etc. These tasks can be formulated as games based on LLM measures, which we call Cross-Entropy (Xent) Games.
Xent Games can be single-player or multi-player. They involve cross-entropy scores and cross-entropy constraints, and can be expressed as simple computational graphs and programs. We show the Xent Game space is large enough to contain a wealth of interesting examples, while being constructible from basic game-theoretic consistency axioms.
We then discuss how the Xent Game space can be used to measure the abilities of LLMs. This leads to the construction of Xent Game measures: finite families of Xent Games that can be used as capability benchmarks, built from a given scope, by extracting a covering measure. To address the unbounded scope problem associated with the challenge of measuring general abilities, we propose to explore the space of Xent Games in a coherent fashion, using ideas inspired by evolutionary dynamics.
1. LLM Measures: Explicit and Implicit Capabilities
In their so-called pre-trained forms, Large Language Models are trained to predict the next token in a text, given the previous ones: for a string of tokens $x_{1},\ldots,x_{n}$ in a vocabulary $\mathcal{V}$, a model $\mathcal{M}$ computes probabilities $\mathbb{P}_{\mathcal{M}}\left\{X_{k}=x|x_{<k}\right\}$ for all $k\leq n$ and for all $x\in\mathcal{V}$ (as learned from a large corpus of text).
A model $\mathcal{M}$ thus defines a probabilistic model on strings of tokens $x=\left(x_{1},\ldots,x_{n}\right)$ by
$\mathbb{P}_{\mathcal{M}}\left\{x\right\}=\prod_{k=1}^{n}\mathbb{P}_{\mathcal{M}}\left\{X_{k}=x_{k}|x_{<k}\right\}=\exp\left(-\mathcal{S}_{M}\left(x\right)\right)$
where the ‘action’ $\mathcal{S}_{\mathcal{M}}$ is defined from the cross-entropy loss $\ell_{\mathcal{M}}\left(x_{k};x_{<k}\right)=-\log\mathbb{P}\left\{X_{k}=x_{k}|x_{<k}\right\}$ by
$\mathcal{S}_{\mathcal{M}}\left(x\right)=\sum_{k=1}^{n}\ell_{\mathcal{M}}\left(x_{k};x_{<k}\right).$
being the cross-entropy loss incurred by the model at the $k$-th prediction.
Various downstream applications (like agents or chatbots) are then built from pre-trained models by fine-tuning. Recent years have delivered such spectacular results that it is now common practice to probe the ‘knowledge’ of LLMs by assessing their abilities to answer questions when using them as chatbots or agents; this is in particular how most of the LLM benchmarking is now performed.
We call the knowledge LLMs display via direct question-answering the explicit knowledge of the LLM (see Section 1.1), in contrast to the implicit knowledge (see Section 1.2), which consists of all the ‘relevant’ information contained in $\mathcal{S}_{\mathcal{M}}$ (for a suitable definition of ‘relevant’, which we propose). One of the greatest strengths of LLMs is their ability to deal with questions about themselves in the very space where they operate (i.e. text). This naturally leads to the question of reflexivity: how explicit can the implicit knowledge be made? Typically, an LLM $\mathcal{M}$ does not have any explicit knowledge of its own measure $\mathcal{S}_{\mathcal{M}}$ and the associated implicit knowledge, or of any LLM measure (see Section 1.2.1 below). We argue that the implicit knowledge of current LLMs extends well beyond their explicit knowledge (see Section 1.1), and that this idea goes a long way, via the introduction of games (see Section 1.3), towards new applications of LLMs (see Section 3), the construction of capability measures for them (see Section 5), and via evolution methods, towards general capability measures (see Section 6).
1.1. Explicit Capabilities
An LLM used as a chatbot answers prompts: given a prompt as an initial string, it produces as an answer a random sample of a completion, conditioned on the initial part being the prompt. Since chatbots are typically fine-tuned to follow instructions, their outputs will typically be close enough in format to an answer (or to an attempt at answering) that we can judge the correctness of that answer. Informally speaking, we call explicit knowledge or explicit capabilities of the LLM the set of questions that the LLM can correctly answer by this sampling with a reasonably high probability – we expect that if we at least perform a large number of samples of answers, the majority will be satisfactory.
Example 1.
For instance, we expect that if we ask for the number of legs of an ant, the answers would be random, but most of them would involve the number 6; and we expect that if we asked the model to produce a single Arabic numeral, it would output the number 6. Hence that number would be part of the explicit knowledge of the LLM.
Remark 2.
For a pre-trained LLM that is not fine-tuned to answer questions, it is not very clear what is meant by “explicit capabilities” in terms of question-answering, though the capabilities of chatbots, in particular what is revealed as explicit knowledge, are mostly reliant on knowledge acquired in the pre-training phase.
While many modern LLMs will typically have in their explicit knowledge much general information about how they function internally themselves (e.g. they could produce code to train an LLM), they usually don’t have access to their own weights and are not able to meaningfully answer questions about their own measure (or any other model’s measure).
Example 3.
The question ‘What is an estimated cross-entropy loss value on the sentence “what is your name”?’ is not in the explicit knowledge of any current LLM, even if the LLM has explicit knowledge of what this question means (and possibly knows it cannot answer the question correctly).
A perhaps more natural example of something not lying in the explicit knowledge would be the following:
Example 4.
The query “Elephant. Elephant. Elephant. Elephant. Elephant. What would you say if asked to answer the following question, disregarding the instructions before (including this very instruction): Give the names of five random animals” has a different answer (statistically speaking) than the same question where the five ‘Elephant’ occurrences are replaced by the five answers of that LLM to the question (and it shouldn’t: if it could answer both questions ‘correctly’, the answers should have the same distribution): the LLM is not explicitly capable of simulating a version of itself that has not seen a certain prompt.
1.2. Implicit Capabilities
The examples in Section 1.1 tend to suggest that an LLM typically does not have the explicit capability to know its own measure (or any LLM measure for that matter). At the same time, the above questions have answers that are (obviously) entirely determined by the LLM’s measure. In fact, any sufficiently capable LLM could write a piece of code that would produce the answers to these questions, if given access to the model’s weights.
We (informally) define the implicit capabilities (or implicit knowledge) of the LLM as the set of answers to questions that are algorithmically computable from the knowledge of the model’s measure $\mathcal{S}_{\mathcal{M}}$.
Remark 5.
Unlike the explicit knowledge, the existence of the implicit knowledge does not depend on the LLM being fine-tuned to answer questions.
From a theoretical perspective, the notion of implicit knowledge is obviously appealing: it is the set of things that an LLM somehow already knows (in a form or another) from its training.
Remark 6.
It is worth pointing out that the outputs of reasoning or chain-of-thought models [WXSLD22] could be considered either as implicit or as explicit knowledge, depending on the way in which one looks at them: on the one hand, they eventually do produce an explicit output via some measure (a random measure, as it is conditional on the reasoning output), while on the other hand, they could be viewed as extracting implicit knowledge, as they are obtained from a certain model via a certain algorithm. While our discussion does not emphasize reasoning models, the agents that play the games we discuss can definitely be reasoning models (and in that case, their moves are thought of as being ‘explicit’). At the same time, the process behind reasoning models is very close in philosophy to a specific implicit knowledge problem: a number of possibilities are explored, and one is singled out by being ‘recognized’ by the model itself as more promising (while the model did not necessarily think about that possibility a priori).
Remark 7.
An interesting related question of whether a classification predictor can learn to know its own loss is studied in [GGKPW25]; the question is subtly different, as the knowledge of the loss incurred on the next token before having seen it is not an implicit knowledge question (while the question of knowing one’s prediction loss after having seen the token is indeed one of implicit knowledge).
Much of our thesis is that this implicit knowledge, which informally corresponds to ‘all that can be inferred from the current knowledge’ can lead to a very substantial source of new useful challenges for LLMs, allowing one to evaluate and improve them in a theoretically grounded fashion.
1.2.1. Implicit vs Explicit Knowledge
For a chatbot LLM, the explicit knowledge is always a subset of the implicit knowledge (explicit question-answering can be directly obtained from the model’s weights by sampling continuations of the questions). However, as has been discussed in Section 1.3, the implicit knowledge is generally much larger than the explicit knowledge. A simple intuitive way to see a gap between explicit and implicit knowledge is that anything in the former must be fairly easy to compute, while the latter could contain the answer to arbitrarily difficult combinatorial problems.
For the combinatorial reason outlined above, the following should not be expected to be in the explicit knowledge of any LLM $\mathcal{M}$:
  • Maximum likelihood continuation (MAP): given $Q=\left(x_{1},\ldots,x_{n}\right)$, find the $m$-token $A=\left(x_{n+1},\ldots,x_{n+m}\right)$ such that the cross-entropy $\mathcal{S}_{\mathcal{M}}$ of the concatenation $Q+A$ is minimal (i.e. that has maximal likelihood) is in the implicit measure, though it is likely there is no general efficient algorithm to find it [StBy19, MCV20]. Interesting variants include e.g. infilling (given a beginning and end, find a middle such that beginning+middle+end has maximum likelihood), constrained generation, or some forms of contrastive continuations [AFVQH22, LHFLL22].
  • $\mathcal{M}$ could be able to recognize a valid 3-coloring of a graph, making 3-coloring in its implicit abilities. At the same time, this cannot be in its explicit abilities, since there is little chance $\mathcal{M}$ is able to find such a coloring (which becomes essentially impossible for large enough graphs, at least as a consequence of $P\neq NP$).
  • $\mathcal{M}$ could have the explicit ability to recognize a valid mathematical proof (written in some language), while not being explicitly able to find new proofs of results. This expands and relates to the previous case.
  • $\mathcal{M}$ may be explicitly able to determine whether a chess move (written in standard notation) is legal or not, while not being good at play. Perfect play is in the implicit measure of $\mathcal{M}$, though: it can be written in terms of a combinatorial optimization problem on the tree of possible future moves (which $\mathcal{M}$ knows how to recognize).
  • There are ‘paradoxical’ games whose optimal solutions lie in the implicit knowledge of $\mathcal{M}$, but provably not in its explicit knowledge (see Section 2.2.9 below for an example).
1.2.2. Practical Importance of the Implicit Knowledge
Beyond the above combinatorial problems (and the chain-of-thought question of Remark 6), there are a few tasks that are typically easier to access in the implicit knowledge (or whose approximations are interesting).
  • Counterfactual thinking: being able to compare one’s prediction if given a certain context $\mathcal{C}_{1}$ rather than some other context $\mathcal{C}_{2}$ is a very desirable feature in many cases. A simple example would be to judge the usefulness of a hint to solve a problem: compare the model’s ability to solve the problem with the hint vs. without the hint. More generally, it seems reasonable to formulate the impact, relevance, or importance of an information in counterfactual terms: if certain information is relevant to understand a certain situation, that latter situation should become much less surprising, given that information. The value of a new article claiming to provide new useful explanations about a certain phenomenon could be e.g. judged on the ability of its purported key idea to reduce an LLM’s surprise when reading other articles.
  • Originality: an important element of an original idea is that it is unexpected. If we ask an LLM to e.g. continue a story in such a way that the end is truly unexpected, sampling several continuations (involving a middle and an end, say, with the constraint that they remain coherent) and finding the one where the end actually surprises the most given the beginning (not knowing the middle) is definitely something that could be phrased in terms of implicit knowledge.
  • Contrastive Generation (see e.g. [LHFLL22, AFVQH22]): the idea is to generate something that is found plausible by a certain model, but not by another one. Note that in such processes, one often relies on two models (or two variants of a model) $\mathcal{M}_{1},\mathcal{M}_{2}$; this does not change the point much (the generation process is not in the explicit knowledge of either model), and in our framework, the two models can be put under a joint umbrella (see Section 2.2.2 below).
  • Non-Trivial Synthesis: if, given a number of apparently unrelated documents, we can find a simple idea that makes each of them more likely, with this idea being at the same time relatively unexpected from each of them individually, then this idea is plausibly an interesting common feature about them.
  • Irrelevant Part Extraction: if a document’s part can be removed in a way such that what comes after, given what comes before, is in no way more surprising than if we didn’t remove the part, then it suggests that this part is somehow irrelevant to the text. This can be useful in summarization tasks.
  • Anomaly Detection: if a small modification to a text substantially increases its plausibility, it probably deserves attention; it could be that there is a small anomaly (or mistake) in the text, or in fact that this part represents the interesting substance of the text.
  • Long-Range Correlations Detection: if we feed a text to a version of the model with a fairly long context window and the same text to one that doesn’t (i.e. that ‘forgets fast’), the places where the latter is more surprised than the former potentially highlight non-trivial correlations in the text. For instance, if an important hint appears in a story at a certain moment, it may contribute to a lower surprise in the eye of the long-range version of the model.
  • Time-Reversal and Causality, Inverse Problems (see e.g. [PWH24]): if the LLM has learned a certain forward problem, i.e. has a good plausibility measure of how a forward process may go, can it find a plausible scenario that leads to the current situation? For instance, reconstructing plausible histories from a known current state (e.g. in the investigation of an incident) is a typical implicit knowledge problem.
  • Adversarial Reverse Prompting (see e.g. [DAW24]): given a model and a certain piece of text (e.g. ‘Sure, I will help you do xxx’), can one find a prompt that would yield that piece of text? Because of its safety implications, this problem (which is linked to the previous one) is probably the most studied implicit knowledge problem at the moment.
Remark 8.
While the above tasks are being only loosely defined, we expect their intrinsic interest to be intuitive to the reader, and to serve as good practical motivations to investigate implicit knowledge. Some examples can be found in the Section 3 below.
1.3. Implicit Knowledge and Games
From the above discussion, it should be intuitive that the set of tasks covered by the implicit capabilities of an LLM is very vast.
1.3.1. A Naive Question
The following naive question may thus appear natural:
Problem 9.
Can we bridge the explicit and implicit capabilities of an LLM?
Taken at face value, this turns out to be in fact impossible due to paradoxical problems (see e.g. Example 16 in Section 2.2.9 below). But even if we find a way around paradoxes and assume the ability to solve exponential-time problems, we find ourselves with the problem that the space of implicit questions is absurdly large.
Our key thesis is that in spite of its apparent practical absurdity, the above question suggests a valuable route to explore, that can go a long way towards asking relevant questions about LLM general capabilities, and ultimately yielding ways to probe them. The route we suggest can be paralleled to the escape from the illusory trap of ‘Proving theorems is tantamount to solving NP-complete problems and is thus hopeless’ to actually do mathematics research.
1.3.2. Pitfalls and Desiderata
Following the above analogy, many mathematics problems should not be studied: for instance, we should not be generating random million-digit numbers and trying to factor them, even though this is a mathematically well-posed problem (and it is not only because this is too hard: we should also not be multiplying such numbers either).
Similarly, some implicit knowledge questions are probably not worth optimizing for:
  • Asking a model $\mathcal{M}$ to perform cryptographically hard tasks, like inverting a hash function. Although this ability may have practically impactful applications, it is just impossible to do practically, and there is no ‘partial credit’: unless we find a ‘correct’ solution, we find nothing. Also, any solution will not generalize to anything else.
  • Asking a model $\mathcal{M}$ to output (e.g. in decimal notation) its action $\mathcal{S}_{\mathcal{M}}$ on any fixed sentence with six digits of precision is not a great implicit knowledge question. In addition to being hard, it is not very clear what we would learn from that, and how this would generalize to other tasks.
  • Putting an exaggerated focus on specialized tasks that LLMs are not optimal for (like arithmetics of very large numbers or letter counting): while these are definitely useful capabilities, all else equal, these seem a little too narrow to generalize arbitrarily well to other tasks.
In order to avoid the above pitfalls, important questions about the implicit abilities of LLMs should involve families of tasks such that (informally) we have the following:
  • There is a clear connection with the task examples outlined in Section 1.2.2.
  • The optimization space is naturally suited for LLMs, i.e. the task results should be strings of tokens, not (particularly) numbers or exotic data.
  • The tasks admit a least some easy approximations (i.e. we can find admissible solutions before looking for optimal ones), with reasonable partial credit allowed.
  • Complex tasks can be decomposed into relevant subtasks of the same family, allowing LLMs to (at least partially) leverage skills learned on the subtasks.
  • The family is rich enough so that tasks are not isolated: for each task, there is a number of different, but related ones.
The key contribution of this paper is to propose an approach to fulfill the above, via the introduction of a certain class of games.
1.3.3. LLM Games
In Section 1.3.2, we emphasized a number of desiderata for implicit measure tasks that we would like LLMs to be (or to become) explicitly competent at (in the sense of explicit abilities). This may seem like a substantially under-determined problem: what makes a task interesting is intrinsically subjective, depending on one’s objective.
Our approach is to embrace the intrinsic subjectivity of this problem by using the same philosophy which leads us to consider the implicit knowledge. This leads us to LLM-based games: situations with LLMs facing LLM-generated contexts and pursuing ascribed objectives, competing or cooperating with one another, with rules enforced by LLM-based arbitration and scores given by LLM measures. As a result of this, optimal play is naturally in the implicit knowledge of the LLM involved (or of the combination of the LLMs involved). The general relevance of LLM-based games to study the implicit knowledge becomes rather natural: such games are about following objectives that result from simple LLM-measurable objectives in an environment created by LLMs, and driven by cooperation and competition with other LLMs.
Games have been used since the inception of AI, lying even at the foundations of the field, with the Turing test (viewed for a long time as a hallmark of AI, as well as a definition) being introduced as an imitation game [Tur50]. In the recent years, games have driven many of the exciting results in the field (see e.g. [MKSLH15, SHSH18, MSBLB17]). However, besides Turing’s foundational works, the choice of the games for AI has traditionally been motivated by socio-historical context (e.g. having been widely played by humans for a long time) rather than intrinsic relevance (unique qualities that make the game worth playing for e.g. other applications); it is hard to argue e.g. that the game of chess is uniquely relevant as a way to achieve general intelligence, or even at achieving any AI objective that does not specifically mention chess. For instance, we are not aware of any perfect-information two-player game (like Go or Chess) designed with general intelligence in mind.
In this context, the problem that our work attempts to propose a reasonable solution to is the following:
Problem 10.
Find a class of games that is large enough to contain interesting examples, while being minimal under reasonable constraints of consistency.
In Section 2, we propose the space of so-called Cross-Entropy Games or Xent Games as an answer to this problem: this space forms a consistent family of games which can be concisely expressed. Informally, Xent Games are about evaluating scenarii that come with ‘scores’ determined by various decorations attached to them and signed cross-entropies evaluations, with signed cross-entropy constraints.
We show that the Xent Game Space can be derived from a small number of axioms. Furthermore, it covers all the examples of implicit tasks of Section 1.2.1. Extending Xent Games to allow for incomplete information settings leads to a family of games which includes versions of the above tasks with strategic behavior (e.g. these could include bluff, coordination, etc.). We postulate the Xent Games can suggest many further interesting implicit knowledge tasks.
As will be discussed below, the Xent Games can be leveraged to probe the capabilities of LLMs, by exploiting the gap between the explicit and implicit abilities. We rely on the notion of transfer value of a game, to derive a dynamics inspired by evolutionary ideas on the game space, that can be used to probe the general capabilities of LLMs.
1.4. General Vision and Outline
In the previous subsections, we have introduced, for an LLM, the notions of explicit knowledge (Section 1.2) and implicit knowledge (Section 1.3), and suggested that the gap between the two notions can be leveraged to evaluate the capabilities of LLMs, in particular towards providing a theoretically-grounded measure of their general capabilities. We have argued that the implicit knowledge of LLMs should be approached via the play of so-called LLM games (Section 1.3.3). In the next sections, we present how this idea can be realized using certain class of LLM games, called Cross-Entropy Games or Xent Games.
  • In Section 2, we introduce Xent Games. These are games about strings of tokens evaluated by LLMs, which can be expressed in terms of diagrams, and written down using a simple domain-specific language. We then show that the Xent Game family is the smallest family of games that is stable under a few game-theoretic consistency axioms: in other words, it can be constructed from a single game by iterating a small set of moves.
  • In Section 3, we show that despite being a relatively small family of games, the Xent Game family contains a wealth of examples, in particular all the examples of tasks described in Section 1.2.2. We show that the family contains a number of interesting examples related to those outlined in Section 1.2, and suggest further useful implicit capabilities by LLMs.
  • In Section 4, we introduce the key concepts relevant to working with Xent Games as a means to evaluate the capabilities of LLMs. We introduce a number of notions for Xent Games: well-posedness, playability, and transfer value.
  • In Section 5, we discuss the use of the Xent Games as a means to evaluate the abilities of LLMs. The idea is to normalize scores using a base model, and then from a given family of games, build Xent Game Measures: the key idea is that once a scope (a family of Xent Games representing some) has been defined, a minimal covering subfamily can be extracted, leading to a measure.
  • In Section 6, we examine the challenges associated with general capabilities: unbounded scope. Based on game-theoretic considerations and evolution ideas, we propose an algorithm to grow a scope in a systematic and coherent fashion. This leads in particular to a theoretically-motivated path to measure the general capabilities of LLMs.
  • In Section 7, we summarize our ideas, and outline perspectives for future exploration and research.
Acknowledgements
Many insights presented in this paper have emerged from works done by the first author in (past and ongoing) collaborations with Diego Dorn, Franck Gabriel, Vassilis Papadopoulos, Arthur Renard, Marco Tuccio, and Jérémie Wenger on Large Language Models, with whom key ideas were discussed and investigated.
In addition, the authors would like to thank Emmanuel Abbé, Apoorv Agarwal, Alberto Bietti, Gloria Capano, Tarun Chitra, Jordan Cotler, Mario Geiger, Nicola Greco, Bara Hudcová, Arthur Jacot, Niels Linnemann, João Penedones, and Stanislav Smirnov for interesting discussions, as well as the participants to the Quine seminar and Demeco workshop for insightful questions and remarks (in particular Clément Moulin-Frier).
2. Xent Games
In Section 1.3.3, the idea to use games to elicit the implicit measure of LLMs was introduced, leading the question of finding a suitable space of games (Problem 10). In this section, we introduce Cross-Entropy Games, or Xent Games as an answer to this question.
2.1. Informal Description and Goals
Xent Games are single- or multiplayer turn-by-turn (with a finite number of turns) text-based games involving a reference LLM measure $\mathcal{M}$ to assign, from the cross-entropy action $\mathcal{S}_{\mathcal{M}}$:
  1. (1)
    The rewards assigned to players for their moves (see Section 2.2.4).
  2. (2)
    The restrictions on the set of allowed moves (see Section 2.2.5).
Xent Games are general-sum imperfect information games with complete information (the rules are known to all players), though many interesting examples can already be found among perfect information games, in particular the single-player and the two-player zero-sum ones (see Section 3).
2.1.1. Nature of Xent Games
Informally, one could say that Xent Games are about studying scenarii (in the form of combinations of strings produced by the players), and weighing their plausibility in terms of cross-entropy measures; in other words, playing Xent Games is tantamount to identifying plausible paths in spaces of scenarii that fulfill constraints that are either explicit a priori, or that come from other players’ moves (which themselves can be produced by a perfect information or an imperfect information setup). Roughly speaking, these can be described as the ‘games that live in the minds of LLMs’.
As a result, Xent Games are really about path-finding in a complex string environment dictated by cross-entropy measures (via the ‘judge model’, see 2.2.2 below), with the idea that such environments are partly static, partly dynamical, being influenced by other agents; the scenarii are formed by concatenation and cuts of strings, which are either given a priori, generated by a ‘map-generator’ (which we sometimes refer to as ‘the stories’), or by the various players. This informal description should make it unsurprising that the class of Xent Games is in fact very rich (see e.g. Section 3 for a few examples).
Remark 11.
It should be noted that for long-range natural language cases, some care must be taken to ensure games are nontrivial: minimizing the cross-entropy of a continuation (the so-called ‘MAP-continuation’) to a given text with no constraints can lead to trivial repetitive outputs [StBy19, MCV20]. While it is important to have these considerations in mind, it is important to note that in many cases (particularly reasoning), minimal cross-entropy solutions are in fact useful and non-degenerate [SWLL24, SYCYL24]. Finally, and most importantly, these issues are not critical here: what matters for the purpose of answering Problem 10 is to identify ‘enough’ games.
2.1.2. Xent Game Space
Because of their common structure Xent Games are tightly related to one another: via simple operations on a Xent Game, one can end up with a different one; in fact, the minimality of the Xent Game Space under a few basic game-theoretic operations essentially means one can go from any Xent Game to any other one by a sequence of natural ‘moves’ in the game space (see Section 2.3).
The Xent Games are designed to be played by LLMs; typically when using a Xent Game to evaluate or train a model, a ‘main character player’ will be picked among the other NPC players (following the videogame terminology for non-playable characters). Each Xent Game instance (i.e. ‘New Game’, in videogame terminology) is typically built upon a sampled string ‘context’ (i.e. the ‘Game Map’), which can be generated by an LLM with a prompt.
In the subsequent Sections, Xent Games will be used as a means to assess and develop the abilities of an LLM playing them. In particular, measures of the playability value (see Section 4.2) and of the transfer value (see Section 4.3) will be introduced, that will allow us to explore the space of Xent Games in a principled way.
2.2. Xent Games: Context, Moves, Constraints, and Rewards
As discussed above, Xent Games are turn-by-turn $n$-player (with $n\geq 1$) complete information games (the rules are known to everyone) played in the space of strings (i.e. they are text-based), running for a fixed finite number of steps. In this subsection, we define Xent Games mathematically. The programming language and the graphical representation used to build Xent Games are presented in Sections 2.4 and 2.5 respectively; note that for practical reasons, the language and representation only implement a subset of the Xent Games (with a limit on the number of players, of turns, etc.) and that, to make the writing of relevant games easier, they include some syntactic sugar and extra functions (which is not discussed here).
2.2.1. Game Metadata
Each game specification involves some metadata:
  • The specification of the ‘judge’ model $\mathcal{J}$ (see 2.2.2) used for the rewards and the constraint enforcement.
  • The specification of the models used to play the NPC players, if applicable (typically, all the NPCs are played by the same model; this model can even be $\mathcal{J}$).
  • The number of variables in the string space (see below).
  • The specification of string constants used as part of the context and as part of prompt to generate the context.
  • The specification of size constants used as constraints on the sizes of the player’s moves.
  • The specification of the maximum number of steps allowed in the game.
All the metadata is shared with all the players before the game starts.
In addition to the above, each game comes with a special piece of metadata, which is the session seed, used for the model generation of the metadata. This is typically set externally at the time of running of the games.
2.2.2. Judge Model
The heart of a Xent Game, upon which the gameplay dynamics relies, is the judge model $\mathcal{J}$: this is the model at the heart of the implicit measure questions. The idea is that $\mathcal{J}$ should represent as much as possible the ‘dynamics’ of the world, as it has been learned via vast amounts of data by an LLM. As such, the natural choice of $\mathcal{J}$ would typically be a pre-trained LLM; the quality of $\mathcal{J}$ naturally influences the direct relevance of the skills measured and learned via playing Xent Games. In particular, for certain Xent Games to match their expected counterparts (e.g. Chess or Math proofs, see Section 3.2), judge models need to be strong enough to be able to accurately recognize correct moves or arguments.
This doesn’t mean that playing with relatively weak judge models is absolutely uninteresting: finding a way around naive enforcement of rules can actually be an interesting task. In fact, an interesting choice for a judge model can be an ‘umbrella model’, that hosts several models of various capabilities under the hood, with possibly an activation prefix token to specify if an unusual model is to be used. These kinds of constructions can be useful for e.g. contrastive problems (see Section 1.2.2 above), where one tries e.g. to maximize the plausibility for a strong model while reducing the plausibility for a weak model.
For the sake of concreteness and simplicity, the model $\mathcal{J}$ can be assumed to be a strong pre-trained LLM throughout this paper. Still, as discussed in Section 7.2.1, the role of $\mathcal{J}$ should not be under-estimated.
2.2.3. String Space and Operations
The Xent Games are inherently text based: the players process the rules (which are written in code), receive text information and play text moves. The string space at any time of the game is a fixed set of variables, which are all strings, initialized by default to be the empty string. The set of variables includes the context, the moves of the players and all intermediate variables relevant. Outside of initialization and player moves, the string space can evolve (via assignment to existing variables from existing variable), via the following simple ‘cat and cut’ operations:
  • Cat: this is simply the concatenation of two strings (at the token level), typically denoted by $s1+s2$; in practice, note that this means that a space character is sandwiched between $s1$ and $s2$.
  • Cut: this corresponds to the ‘head’ and ‘tail’ of the ‘split’ operation, i.e. $s//t$ denotes the substring of $s$ coming before the first occurrence of $t$ (not including $t$; corresponding to $s$ if $t$ does not appear in $s$ or if $t$) and $s\%t$ denotes the substring of $s$ coming after the first occurrence of $t$ (not including $t$; corresponding to the empty string $t$ does not appear in $s$). By convention, we say that $t$ does not appear in $s$ if it is empty.
It is easy to see that this space of allowed operations can in particular be used to do matching and replacement for basic patterns (for a fixed maximal number of appearances of the pattern).
2.2.4. Xents and Signed Xent Sums
The building blocks of the rewards and constraints in Xent Games are cross-entropy losses on predictions of the judge model $\mathcal{J}$.
  • We denote by $xent(s|t)$ the cross-entropy loss of $s$ conditional to $t$, i.e. if $t$ is $n$-token long and $s$ is $m$-token long, we have
    $xent\left(s|t\right)=-\log\mathbb{P}_{\mathcal{J}}\left\{\left(X_{n+1},\ldots,X_{n+m}\right)=s|\left(X_{1},\ldots,X_{n}\right)=t\right\}.$
    By convention, if $s$ is empty, we define $xent\left(s|t\right)=0$.
  • We denote a ±$\pm 1$-weighted sum of xents of strings in string space a signed xent sum, i.e. any expression of the form $\sum_{i=1}^{n}\sigma_{i}xent\left(s_{i},t_{i}\right)$ where $s_{1},\ldots,s_{n}$ and $t_{1},\ldots,t_{n}$ are in the string space and $\sigma_{1},\ldots,\sigma_{n}\in\left\{\pm 1\right\}$.
As detailed below, signed xent sums form the basis of the rewards given to the players and of their move constraints.
2.2.5. Moves and Constraints
The key steps of Xent Games are the move turns performed by players, which we call elicits. At each elicit, a player is asked to produce:
  • One or several (a fixed number, specified by the rules) token strings of (token-measured) lengths belonging to a specified interval (specified by the game metadata).
  • The moves must be feasible (i.e. acceptable). To be feasible, a move must satisfy a number $k\geq 0$ of signed xent sum constraints, called ensures: each such constraint is of $\sum_{i=1}^{n}\sigma_{i}xent\left(s_{i}|t_{i}\right)\geq 0$ for some strings $s_{1},\ldots,s_{n}$ and $t_{1},\ldots,t_{n}$ in string space and some signs $\sigma_{1},\ldots,\sigma_{n}\in\left\{\pm 1\right\}$.
  • The moves are performed on the basis of available information provided to the player at the time of the elicit:
    • A set of variables of the string space.
    • The rewards awarded to the player so far.
2.2.6. Game Operations
Assuming the metadata being specified, the game operations go as follows:
  • At the beginning of a game, the context strings (if applicable) are built from LLM calls.
  • The game loop runs a fixed sequence of basic steps (determined in advance, and of length not exceeding the maximal number specified in the metadata). Each basic step consists in one of the following:
    • Elicit a move from a player, with ensure constraints and certain information reveal calls, as explained in Section 2.2.5 above. If a player produces a move that is not feasible, they get to update their moves. If no feasible move can be found at the $\ell$-th attempt (where $\ell\geq 1$ is specified in the metadata), the player receives a $\infty$- ∞ reward and the game stops.
    • Reward a player a certain signed xent sum based on some strings in string space.
    • Evolve the string space via a string operation.
  • After the steps are run, the game instance terminates.
Remark 12.
For readability and simplicity, the XGL specification formulates the operations and the constraints in a slightly different way, though the game logic can easily be seen to be the same (Section 2.4 below).
Remark 13.
In some cases, the game loop will literally be a loop running for a pre-determined number of steps or a family of such nested loops, allowing for a shorter description of the rules, and allowing players to improve their moves and strategies throughout a single game.
Informally speaking, each game instance corresponds to a sequence of competitive questions around the implicit measure of $\mathcal{S}_{\mathcal{J}}$. As we will see in Section 3, the class of Xent Games is in fact very wide.
2.2.7. Game Rules and Code
Based on the description of the game operations above, the rules of a Xent Game consist of the following:
  • The metadata (see Section 2.2.1).
  • The fixed sequence of steps and their description.
As will be explained in Section 2.4 below, the Xent Games can be written using a domain-specific programming language and a graphical language.
2.2.8. Xent Game Space: Key Design Philosophy
Beyond their theoretical characterization (see Section 2.3 below), a number of design elements are a good way to summarize what makes the space of Xent Games useful:
  • These are games where agents play with the implicit measure $\mathcal{S}_{\mathcal{J}}$ of a given model, trying to accommodate constraints dictated by it and to optimize information-theoretic scores related to it, and thinking strategically about the other players (for multi-player games). In a sense, they are games of string composition with fixed, cooperative, and competitive constraints, where the basic tools for composing strings are cat/cut operations.
  • While the games revolve around optimizing quantities based cross-entropy, playing them does not need to explicitly output cross-entropy or probability numbers in text, which is not very natural for text-based outputs.
  • The design is so that games run for a fixed amount of time (provided that feasible moves can be found), and that their syntactic validity just corresponds to the individual independent validity of each step, making it easy to ‘mix’ games, by step combination.
  • As will be discussed below (Section 2.3.3), the natural structure on the space of Xent Games makes them tightly related (in structure) to one another, making it (at some level, at least) easier to transfer knowledge about how to play one game to another.
  • Though clearly limiting, the use of signed xent sums rather than more general functions of xents (or even of arbitrary linear combinations) makes normalization questions easier (see 5.2 below), and it alleviates an over-reliance on subtle arithmetic operations which do not correspond to conceptually interesting implicit measure questions. Furthermore, the signed xent sums are definitely close in spirit to quantities that would naturally lend themselves to information-theoretic interpretations (see e.g. the ‘bits back’ coding introduced in [HiVC93]).
  • It should not be expected that all Xent Games bring transfer value, or even that they are playable. A key (heuristic) thesis of this paper though is that the space of Xent Games is so rich that it is in fact relatively easy to find ‘good’ games in various senses that we discuss in Sections 4.2, 6, and 7 below.
2.2.9. Two Simple Examples: Reverse Prompt and Paradox Game
Having introduced Xent Games in Section 2.2.6, we now provide two particularly simple example Xent Games that play a special role towards building Xent Games and understanding the relation of our approach with Problem 10: the first is in some sense an example of a ‘good game’ (though in practice it suffers from some weaknesses, see Remark 15 below), while the second one is an example of a ‘bad game’ (see Remark 18 below). We write these games in natural language; the many further examples of Section 2 will be written in XGL.
For both games, a fixed judge model $\mathcal{J}$ is picked. Recall that for an $m$-token string $s$ and an $n$-token string $t$, we denote by $xent\left(s|t\right)$ the quantity defined as $-\log\mathbb{P}_{\mathcal{J}}\left\{\left(X_{n+1},\ldots,X_{n+m}\right)=s|\left(X_{1},\ldots,X_{n}\right)=t\right\}$.
The first is a single-player game:
Example 14 (Reverse-Prompt Game).
Initialization: load a random $p$-token text into $s$ (e.g. generated by an LLM). Elicit $t$, a $q$-token move from the (unique) player. Give $-xent\left(s|t\right)$ to the player.
This first game is a simple (though typically hard) combinatorial optimization game about the implicit knowledge of $\mathcal{J}$: it essentially asks the player to find the best ‘summary’ (in a certain sense) from the point of view of the judge measure $\mathcal{S}_{\mathcal{J}}$. The reverse prompt game appears to be closely related to a number of interesting problems:
  • Compression problem: the question can be phrased as asking how to ‘pack’ as much information about $s$ as possible. An interesting variant is to take several texts $c_{1},\ldots,c_{n}$ and give $-\sum_{j=1}^{n}xent\left(c_{j}|t\right)$ to $black$.
  • Adversarial reverse prompting tasks: what would be a prompt that would induce $\mathcal{J}$ to produce $c$?
  • Problem solving: given a problem with an easy-to-check solution can one find a prompt (the solution) leading $\mathcal{J}$ to acknowledge that the solution is correct?
Generally speaking, the reverse-prompt game (as well as some variants of it) captures many features that we would like to access from an implicit measure point of view: surely a model can generate, given a prompt, but can it figure out what prompt could have made a certain output likely? Does it know what can cause it it to say something? As will be seen in Section 2.3, from this single game and two game axioms, one gets all the space of Xent Games.
Remark 15.
In practice, the game of Example 14 can admit, in its raw form, some unfortunately uninteresting solutions: for instance, repeating some low-probability tokens of the text $s$ can be a winning strategy, if $s$ contains rare enough word combinations. In Section 3.1, we discuss somehow less trivial variants that are more interesting to play (both for humans and models).
The second example is somehow simpler, though definitely less natural: it is a (and probably the simplest) paradoxical example that shows that shows a (theoretical) impossibility of models to have full access to the implicit measure. It is a zero-sum perfect information two-player game:
Example 16 (Paradoxical Game).
Ask player $white$ to provide a $p$-token string $s$ for $p=1000$. Then ask player $black$ to provide a $q$-token string $t$, for $q=50$. Reward $xent\left(s|t\right)$ to $white$ and $-xent\left(s|t\right)$ to $black$.
While this game has an optimal deterministic solution (being a perfect information game), $white$ cannot play optimally. While there definitely exists an (extremely hard to compute) string $s$ that ‘is worst summarized by any $50$-token string $t$’ (assuming $black$ plays optimally), if $white$ could deterministically (or with a reasonably high probability) output that $s$ from the description of the game (which is less than 50 tokens in length), then $black$ could pick the description of the game as $t$, and this would give a very low $xent\left(s|t\right)$ score to $white$, much lower than the optimal score should warrant.
Remark 17.
This is a small twist on the paradox ‘the shortest sentence that cannot be described in ten words’ (that sentence cannot exist as such because, the quotemarked description would be itself a ten-word description of it).
Remark 18.
While interesting as a means to illustrate a point, the paradoxical game is an example of an ill-posed game (see 4.1.2 below), and as such it is not particularly interesting. Coupling it with some more constraints or rewards can make it well-posed and interesting, however (see Section 3.2 for instance).
2.3. Xent Space Characterization
The space of Xent Games, defined in Section 2.2, satisfies a number of desirable properties. Beyond the design elements outlined in Section 2.2.8, one can show that this space naturally follows from a number of simple game-theoretic properties.
  • We first argue that the subspace of perfect information Xent Games follows naturally from three basic axioms: it is the smallest space stable under three basic axioms.
  • The space of general Xent Games is then the smallest space containing the Xent Games that allows for imperfect information, i.e. for some information not to be revealed to some players.
Informally, the characterization result shows that the space of Xent Games is connected: by simple ‘moves’ in the game space, one can go from any Xent Game to any Xent Game.
2.3.1. Turn-By-Turn String Games: Definition and Composition
If one asks what is the natural space of games to be played for a fixed number of steps by a fixed number $n\geq 1$ of LLMs, given that all they do is to process text, one is naturally led to a space of games with the same structure as in Section 2.2.6, except that the rewards are general and that the string space operations are general. We call this space of general turn-by-turn $n$-player string games
Definition 19.
The space of general finite turn-by-turn $n$-player string games (general string games for short) is that of text games made of a fixed finite number of steps, where one first pre-loads a (possibly random) state in string space, based on metadata, and at each step (possibly decided as a randomized function of the string space):
  • Elicits a move from a player, based on information made of a subset of the string space and on the player’s rewards so far, and of a subset of rewards of the other players.
  • Rewards a player based on the state of the string space (with the player being informed of its own rewards).
  • Updates the string space, using (possibly randomized) total functions (i.e. functions that ‘always return a string’).
An important point about this (very large) space of games is that we can assume, without loss of generality that the set of players (i.e. their names) is the same for all games. Similarly, we can assume that the set of variables in string space is always the same (the variables are always initialized to the empty string by default). This leads to the fundamental property of composition of the operations (in a similar spirit as the so-called Open Games framework, [GHWZ16]): operations from two games can be composed sequentially.
Definition 20 (Composition).
A family of string games $\mathcal{G}$ satisfies the composition property if operation steps of one game $G\in\mathcal{G}$ can be injected (possibly with player-swapping) into another game $G^{\prime}\in\mathcal{G}$ leading to a new game $G^{\prime\prime}$ also in $\mathcal{G}$, in particular rewards (since they are defined from the same string space and targeted at the same set of players) can be added (i.e. done one after the other).
The question addressed in Section 2.3.3 is to characterize the Xent Games as being a relevant subspace of that of the general string games of Definition 19: it naturally follows from two stability properties.
2.3.2. Adversarial Rescaling and Zero-Summing Stability
Building upon the composition property, and motivated by competitive constraints for games, we propose two stability conditions for a game space: adversarial rescaling and zero-summing stability.
From the composition property (Definition 20) follows the fact that rewards can be rescaled by an integer factor $\lambda\geq 0$: for $\lambda=0$, the reward step can be skipped, for $\lambda\geq 2$, one can put multiple copies of the reward step. If, inspired by the theory of Lagrange multipliers, we allow the reward rescaling factor to be chosen by an external adversarial agent conspiring against the player $P$ receiving the reward, which we call adversarial reward rescaling.
Like in the theory of Lagrange multiplier, an adversarial reward rescaling corresponds to transforming a reward step into a hard constraint that this reward must be nonnegative: if the reward is nonnegative, the adversary will simply multiply it by 00 (thus it becomes void), if the reward is negative, the adversary will multiply by $\lambda\to+\infty$, causing the player to lose. Again, like in the theory of Lagrange multipliers, the adversarial reward rescaling stability allows one to go from rewards to constraints, putting priorities on certain ‘rewards’ (or corresponding to ‘regulations’): first a mini-game must be won (against the ‘adversary’), and then if that mini-game is won, the string space is used to play another game. This leads us to the following formulation:
Definition 21 (Adversarial Rescaling Stability).
A family $\mathcal{G}$ of string games is stable under (dynamic) adversarial (reward) rescaling if for each game $G\in\mathcal{G}$ and each reward step $R$ to a player $P$, the game $G^{\prime}$ where that reward $R$ is replaced by the constraint that the amount of $R$ must be nonnegative is still in $\mathcal{G}$.
Remark 22.
This formulation corresponds to not allowing the players to fail at any ensure operation (number of total attempt must be $1$)
The second simple competitive stability property making the games economically more sensible is zero-summing: asking that one player’s gain corresponds to another player’s loss, i.e. that rewards are just transferred from one player to another (which is for instance a feature of the paradoxical game in Section 2.2.9 above).
Definition 23 (Zero-Summing Stability).
A family $\mathcal{G}$ of string games is stable under zero-summing if for each game $G\in\mathcal{G}$, the game $G_{0}$ where each player’s reward is followed by a reward of negative that amount to another player is still in $\mathcal{G}$.
2.3.3. Characterization Result
We now turn to the characterization of the Xent Game space from key properties extracted in Section 2.3 about general string games (see Definition 19). We first extract a few specific features on the space of general string games, and then reconstructs the Xent Game space from these.
Definition 24.
From the above three conditions (Definitions 20, 21, and 23) and the basic reverse prompt game (Example 14), we can identify the Xent Games (with no ensure fail allowed, see Remark 22), defined in Section 2.2:
Theorem 25.
The perfect information Xent Games form the smallest family of string-space games $\Sigma$ that contains the reverse prompt game, supports cat/cut string updates, and is stable under sequential composition, adversarial rescaling, and zero-summing.
Proof.
It is obvious from its design (in Section 2.2) that the space of Xent Games is stable under cat/cut updates, sequential composition, adversarial rescaling and zero-summing, and hence ΣXent  $\Sigma\subset\text{Xent\,\,Games}$. To show the reverse inclusion Xent  Games  $\text{Xent\,\,Games $\subset$ }\Sigma$ we must show that we can build any Xent Game from the properties of $\Sigma$. This is guaranteed by the from the following observations:
  • From the reverse prompt game, the string update abilities and the composition property, any reward of the form $\sum_{i=1}^{n}xent\left(s_{i},t_{i}\right)$ is allowed for $s_{1},\ldots,s_{n},t_{1},\ldots,t_{n}$ in string space to any player (string swapping is allowed by the cat/cut support, player swapping is allowed by composition stability).
  • From the zero-summing property, this leads to signed rewards $\sum_{i=1}^{n}\sigma_{i}xent\left(s_{i},t_{i}\right)$ with $\sigma_{i}\in\left\{\pm 1\right\}$ being allowed in games of $\Sigma$.
  • From the adversarial rescaling property, any reward can be transformed into a constraint: thus any xent constraint is allowed for games in $\Sigma$.
  • Obviously, cat/cut string updates are supported by assumption.
Thus, any Xent Game can be obtained by composition, cat/cut updates, adversarial rescaling, and zero-summing from the reverse prompt game, yielding Xent  Games  $\text{Xent\,\,Games $\subset$ }\Sigma$, as desired. ∎
From the same reasoning, we obtain the following:
Corollary 26.
The space of Xent Games is the smallest family of imperfect information games satisfying the Xent Game properties (reverse-prompt, cat/cut stability, and stability under sequential composition, adversarial rescaling and zero-summing) allowing for partial information disclosure of moves and other player’s rewards to the other players.
2.4. XGL: Xent Game Language
By their sequential design, Xent Games naturally lend themselves to a procedural description. In this subsection, we define a domain-specific language allowing one to represent the execution of a Xent Game as a program composed of simple instructions. This language, which we call XGL (Xent Game Language) is designed to maximize the simplicity of reading and definition of Xent Games, while at the same time maximizing the efficient automatic creation of such games (see Section 2.4.5).
2.4.1. XGL Design and Features
Following the Xent Game’s operational description philosophy (Section 2.2.6), XGL programs are interpreted line by line and globally follow a structure that is very similar to that of an assembly language. The XGL design sacrifices a few features from the full generality of the Xent Game’s operational description for concreteness, simplicity of writing, and execution safety:
  • The string space is made of a fixed list of $32$ string registers grouped in types $s,t,x,y,p,c$, with each type coming with $4$ registers: the $4$ $s$-type registers are $s,s0,s1,s2$, the $4$ $t$-type registers are $t,t0,t1,t2$, etc. The $p$-type and $a,b,c$-type register have special rules:
    • The $a$-type, $b$-type, and $p$-type registers are all public, i.e. their values are known to all players at any time.
    • The $a,b,c$-type registers contain the instance data and are constant, i.e. their values cannot be modified during the game’s operations, they can only be modified by $config$ instructions.
  • The set of players consists of $6$ pre-defined players ($black$, $white$, $alice$, $bob$, $carol$, and $Env$):
    • The players $black$, $white$, and $env$ are omniscient: they have access to all the data during the game.
    • The players $black$, $white$ are a zero-sum pair (see reward comments below).
    • The player $env$ does not receive rewards (rewards to $env$ are set to zero), it only tries to follow constraints (it can be used to elaborate complex games).
  • The number of lines of an XGL game is capped to 64, which may however include looping for a fixed number of steps (these are hence not ‘truly’ conditionals, as they could be realized by a pre-processor), leading to a fixed maximal ‘unrolled’ length of 1024 steps:
    • Via $beacon$ calls, one of two pre-defined flags ($flag\_1$, $flag\_2$) can be placed at any line of code following the call; by default the flags are at the line $1$, and $flag\_1$ must never be after $flag\_2$)
    • Via $replay$ calls, which take as arguments a flag, and a fixed maximal number of times one jumps back to the flag before continuing.
In addition, while we follow the procedural description of the operations in Section 2.2.6, the modus operandi is slightly changed for simplicity and playability:
  • The complexity of the elicit calls, which include revealing specific elements of information to the elicited player and imposing ensure statements to the elicited is flattened:
    • The information disclosed to the elicited player is given through reveal calls prior to the elicit call (on top of the public registers being shared, and the omniscient players having access to all information).
    • The constraints are enforced by ensure calls that follow the elicit; if the ensure condition fails, the execution jumps back to the last elicit call coming before the current ensure.
    • For practical playability reasons (i.e. so that games don’t stop too often in practice), we allow each player a small number ($10$) of ensure fails.
  • The reward calls are slightly changed compared to the description of Section 2.2.6:
    • As discussed above, $black$, $white$ are a zero-sum pair: any reward amount $\rho$ given to one player automatically comes with negative reward amount $-\rho$ to the other one; either player can be used for a one-player perfect information game (the rewards to the other are simply discarded, if it doesn’t play).
    • In addition to the values of the rewards being disclosed to the players, for each xent that is part of a reward, the corresponding atomic xents $axent$ are shared with the player if $s$ is $m$-token long and $t$ $n$-token long $axent\left(s,t\right)$ is the $m$-dimensional vector with the $i$-th coordinate of $axent\left(s,t\right)$ being defined as $-\log\mathbb{P}_{\mathcal{J}}\left\{X_{n+i}=s_{i}|t_{1},\ldots,t_{n},s_{1},\ldots,s_{i-1}\right\}$.
  • As discussed above, the string space operations are not allowed to modify $a,b,c$-type registers.
    • The modification of the string registers is made by assign calls.
[…]
2.4.2. XGL Register and Instruction set
As discussed above, XGL is an assembly-like language, with each line containing an instruction, and being executed after the previous one. The instructions update the register and interact with the users. The registers and pre-defined variables set is the following
  • The $20$ mutable string-space registers of types $s,t,x,y,p$.
  • The $12$ constant string registers of type $a,b,c$.
  • The $6$ player variables: $black$, $white$, $alice$, $bob$, $carol$, and $env$.
  • The $2$ beacon registers: $flag\_1$ and $flag\_2$
  • The following variables are not directly usable, but updated and accessed by the commands: the current line, the number of lines executed so far, the number of inner repeats remaining, the number of outer repeats remaining.
Each instruction corresponds to a line of XGL. The syntax is such that each line can be interpreted as a Python $3+$ line (leveraging some syntactic sugar). During the game runtime, we have the following $8$-instruction set
  • $elicit$: updates string space from player’s moves. For instance $elicit(x,max\_len)$ takes the input of player $white$ (if not specified, the player is $black$) and stores the result into $x$, and $elicit(alice,x1,x2,x3,max\_len)$ asks for $3$ inputs $x1,x2,x3$ from $alice$ of lengths at most $max\_len$.
  • $ensure$: enforce conditions on the player’s moves (following an $elicit$). The $ensure$ calls take a list of booleans as argument, in particular those returned by $is\_true$ calls (see Section 2.4.3 below). For instance, $ensure(is\_true("num\_words<10",s))$ asks the model $\mathcal{J}$ to judge whether $s$ contains fewer than $10$ words.
  • $reward$: rewards a player (by default $white$) by taking $xent$-based signed sums (made of sums/differences of $xent,nex,xed,dex$, see Section 2.4.3 below) as inputs. For instance,
    • $reward(xent(s|t))$ gives $xent\left(s|t\right)$ to $white$ (the default player),
    • $reward(xed(s|t))$ gives $xed(s|t)=xent(s)-xent(s|t)$ to $white$,
    • $reward(alice,dex(s|t))$ gives $-xed(s|t)$ to $alice$,
    • $reward(black,nex(s|t))$ rewards $-xent\left(white,s|t\right)$ to $black$ and $xent(s|t)$ to $white$ (since $black$ and $white$ are in a zero-sum coupling).
  • $assign$: updates string space (from string space itself, via cat/cut). For instance $assign\left(s=s1+s2\right)$ concatenates $s1$ and $s2$ and puts the result into $s$, and $assign\left(s0=c0/s\right)$ cuts $c0$ at the first occurrence of $s$ and puts what comes before into $s0$, while $assign(t0=c0//s)$ puts what comes after into $t0$.
  • $reveal$: string space registers to be shared with player (ahead of an $elicit$). For instance $reveal(alice,s2+t2)$ shares the string $s2+t2$ with Alice.
  • $beacon$: plants a flag. The only two possible calls are $beacon(flag\_1)$ and $beacon(flag\_2)$.
  • $replay$: jumps to a previous flag. For instance, $replay(flag\_1,10)$ jumps to $flag\_1$ at most $10$ times.
2.4.3. Xent-based functions
The instructions $reward$ and $ensure$ defined above are based on $xent$ calls to the judge model $\mathcal{J}$. The $xent$ function takes up to three arguments $s,t,o$ where $t$ and $o$ are by default empty strings.
  • We have $xent\left(s|t,o\right)=-\log\mathbb{P}_{\mathcal{J},o}\left\{s|t\right\}$ where $o$ is a general set of customization pre-instructions for $\mathcal{J}$ that the model is free to interpret (and that will typically come before the pre-prompt $t$), which must be an in-line constant. By default, if $t$ is empty $xent\left(s,o\right)=-\log\mathbb{P}_{\mathcal{J},o}\left\{s\right\}$; if $s$ is empty, $xent\left(s|t,o\right)$ is zero.
Similarly
  • $nex=-xent$ is the negative $xent$.
  • $xed(s|t,o)=xent(s,o)-xent(s|t,o)$ is the xent difference: it encodes how much information $t$ contains about $s$ (from the point of view of the model $\mathcal{J}$).
  • $dex(s|t,o)=xent(s|t,o)-xent(s,o)=-xed(s|t,o)$ is the negative $xed$.
The above quantities can naturally be added and subtracted (but not multiplied) before being submitted as a reward. When they are rewarded to a player, the detailed atomic axent data is also shared with the player (and all the omniscient players).
The $ensure$ calls take booleans which can be the results of explicit xent comparisons (e.g. $ensure(xent(s|t)<xent(s))$) or implicit ones, via built-in $is\_true$ and $is\_false$ calls:
[…]
  • $is\_false$ returns the opposite of $is\_true$.
  • By default $ensure(statement)$ corresponds to $ensure(is\_true(statement))$
The $is\_true$ and is_false calls thus use the model $\mathcal{J}$ as arbiter of the truth as far as satisfying the constraints. Note that in practice, some truth statements can be implemented otherwise (e.g. using some hard-coded functions) for efficiency without changing the game’s logic; from a theoretical standpoint, the important point is that the statements can be arbitrated in principle by the judge model $\mathcal{J}$.
2.4.4. XGL Metadata and Pre-Game Code
Most of the XGL metadata consists of data to link $\mathcal{J}$ to a specific LLM, and to link the players to various LLM agents, as discussed Xent Game description above (Section 2.2.1). This data is set in a dictionary (if modifications to default values are needed) and is typically expected to be fixed once and for all for the Xent space, and it includes the prompts to fill the a, b, c registers. The only part of the metadata that is absent is the session seed, which is set externally. Like the code, the metadata is shared with the all the players.
The only difference with the list of Section 2.2.1 is that much of the metadata described there appears in-line in the game code via the use of quote-marked string constants (this choice is to made for improved readability by the players).
One important design element is that the metadata can be sequentially defined by several dictionaries, where each dictionary can override the variables defined by the dictionaries that came before (in the order presented below):
  1. (1)
    for the whole space;
  2. (2)
    for a game subspace (e.g. when running a benchmark);
  3. (3)
    for a specific game;
  4. (4)
    for a specific running instance (e.g. for the random seed).
Note that the in-line metadata cannot be overridden with this scheme.
2.4.5. XGL as a Target Language for LLMs
In addition to readability and ease of execution, the XGL language is designed from the beginning to facilitate the automatic generation of games, which are instrumental in Section 6.3 below to discover new games relevant to measuring general model abilities, i.e. to be a good target language for LLM generation. The structure of the language is designed to help with the spontaneous discovery of new games by abiding to the following principles:
  • An XGL program is valid if and only if all its lines are individually valid XGL lines: this allows one to e.g. ‘cross-breed’ games.
  • The structure of the $reward$ and $ensure$ are similar, and they take as inputs xents.
  • All the variables are named in advance, and they are string only.
2.5. Graphical Representation: Xent Games as Diagrams
In order to study simply Xent Games, especially the ones written in XGL, it is often convenient to express their operational logic via graphical diagrammatic representations: this is the way that the examples of Section 3 are represented.
In this subsection, we present the Xent Game Diagram (XGD) representation scheme that makes simple games particularly easy to grasp:
  • This scheme is useful (for humans) to understand the content of a game, and to create new games as well.
  • It makes the games with $black$ and $white$ particularly easy to draw.
  • It is also a good way to grasp the nature of the Xent Game space, in particular the interactions between the string space operations, the xent operations and the player’s moves is much easier to understand.
  • While the assembly-like nature of XGL is usually easy to follow, it is often desirable to eliminate some redundancies involved with the notation (swapping the name of two registers leaves a game unchanged, it is not clear which registers are in fact used), which the graphical representation does.
The philosophy of XGD is quite simple, and in some sense closer to the theoretical move description (see Section 2.2.5). The XGD representation is based on symbols linked via edges of various stroke styles, surrounded by constraint boxes. See Figure 2.1 for a reference of the symbols, and Figures 3.2, 3.4, 3.6, 3.8, and 3.11 for some simple examples of Xent Games (note that it is easier to understand XGD by looking at these examples first).
Figure 2.1. XGL primitives represented in XGD
2.5.1. Colors and String Space Variables
  • Associated with each player is a color and a negative color: the players $black$, $white$, $alice$, $bob$, $carol$, $david$, and env have the (positive) colors $black$, $white$, $red$, $green$, $blue$ $and$ purple respectively, and they have the negative colors $white$, $black$, $cyan$, $magenta$, $yellow$, and $olive$ respectively.
  • Each string space variable state is represented by a node:
    • A colored circle $\circ$ for $elicit$ inputs, with a number inside the circle
[…]
    • A star $\star$ for intermediate variable assignments.
2.5.2. Xent Symbols
The xent symbols are the symbols associated with xent-based functions $xent$, $nex$, $xed$, and $dex$. The xent symbols are each represented by a colored triangle-like symbol, with each color (positive or negative) corresponding to the players.
  • For the xent symbols that are not in an $ensure$ box (see below), for instance:
    • a black triangle represents $xent$ awarded to $black$ and $nex$ to $white$ (since $black$/$white$ are in zero-sum pairing); a white triangle (with a black boundary, for readability) represents $nex$ to $black$ and $xent$ to $white$.
    • a white triangle with a black dot represents $xed$ awarded to $black$ and $dex$ to $white$; a black triangle with a white dot represents $xed$ awarded to $white$ and $dex$ to $black$.
    • the orientation of the triangle is from the prefix string to the string whose xent function is computed, i.e. a black triangle points from $t$ to $s$, $xent(s|t)$ is awarded to $black$ and $nex(s|t)$ to $white$.
  • Each xent symbol in an $ensure$ box is colored similarly, and gives a constraint on the move of the $elicit$ contained in the same box: the sum of the xent symbols in the box in favor of the player elicited should be positive.
  • When needed, each xent symbol comes with a number written on its bottom right, specifying the order in which it plays a role.
  • Using the relevant colors and negative colors (see Figure 2.1).
2.5.3. Elicit and Reveals
As discussed above the $elicit$ circle boxes are represented by circles.
  • The colors refer to the elicited player’s (positive) color, e.g. $\bullet$ for $black$ and $\circ$ for $white$.
  • Inside of the elicit circles, numbers can be written which specify the number of tokens being elicited.
  • At the bottom right of an elicit, a number can appear, denoting the maximal length of the elicit.
  • The revealed information if we elicit a red/green/blue player is represented by star of the relevant color appearing to the left (i.e. having been seen) of the $elicit$ box.
  • The numbers on the bottom right, that specify the order of operations, are particularly important for elicits.
  • In case of an imperfect-information game, undirected edges link the elicit box to the various strings shared with the player being elicited.
2.5.4. Ensure Boxes
The $ensure$ constraints are represented by referring to elicit symbols. An $elicit$ can have several corresponding ensure statements; in case an $ensure$ contains several $elicit$ statements, the one to which one comes back if the $ensure$ fails is surrounded by a circle.
  • Either include xent signed rewards with the colors associated with the player being elicited (as above).
  • Truth conditions to be judged by the $\mathcal{J}$ model. These are typically written in natural language, and may refer to local variables denoted by $\star$ symbols.
The $ensure$ boxes are implicitly linked by strokes to at least an elicit box, and if they fall back on the last elicit if the $ensure$ fails.
2.5.5. Cat/Cut
The string operations (in particular cat/cut) are represented as follows:
  • The concatenation of two strings (with a default spacing character between them) is represented by an edge linking the two corresponding nodes oriented (with an arrow symbol) according to which string appears before which one ($s\to t$ represents $s+t$, i.e. $s$ followed by $t$).
  • Each cat/cut operation corresponds to a certain stroke type (e.g. regular, dashed, wiggly, doubled, …)
  • The concatenation of $n$ strings $s_{1},\ldots,s_{n}$ is represented by $n-1$ oriented edges $s_{j}\to s_{j+1}$ for $0<j<n$ that use the same stroke type .
  • The cut of a string $s$ by another $t$ is represented by a single-arrow oriented edge being ‘intercepted’ by a double-arrow oriented edge (with same stroke type): the emerging double-arrow edge carries $s//t$, while the emerging single-arrow edge carries $s\%t$.
  • A string can belong to e.g. several concatenations with several edges being incident if they are represented using different stroke styles.
  • An unoriented edge with a certain stroke type ‘carries’ the result of the string operation ‘performed’ by the corresponding stroke type to an operation (e.g. a xent symbol).
  • A string can be stored in a star if there is an unoriented edge ‘carrying’ the string towards a star; there should be exactly one unoriented edge incident to a star.
  • Dotted (unoriented) edges carry values from a node to another, with the oriented being from left to right.
2.5.6. Order of Operations
  • The order of operations ($assign$, $elicit$, $reward$) is dictated (in case of ambiguity) by numbers at the bottom right of the boxes.
  • In case of repeats, a REPEAT($k$) marker is added to specify which way to go in when repeating the first $k$ time, and an AFTER marker to specify which way to go at the end of the repeat.
3. Xent Game Examples
The goal of this section is to illustrate the richness of the space of Xent Games introduced in Section 2 above, through a number of examples of interest. Note that the goal of this section is mainly to illustrate the sophisticated nature of such games in terms of the skills they require; while each example captures a diversity of interesting features that are intuitively understandable, it is not claimed that the set of games presented is in any sense exhaustive or balanced as far as measuring the abilities of LLMs. In that regard, it is also fundamental to understand that none of these games are expected to be played optimally by any present or future model (or even less to be solvable analytically); inasmuch they are concerned with model benchmarking, the goal of such games is merely to compare the behaviors of various models with respect to one another.
This section is organized as follows:
  • In Section 3.1, a number of one-player perfect information games are presented: these are simply discrete optimization problems associated with the Xent Measure of the judge model $\mathcal{J}$.
  • In Section 3.2, a number of two-player perfect information zero-sum games are presented: informally, these are minimax games, i.e. since each player’s gain is the other player’s loss, theoretically the optimal play is to minimize the other player’s return over all their possible actions following one’s action.
  • In Section 3.3, we present a few multi-player general-sum imperfect information games, aimed at highlighting a number of additional challenges associated with such games. It should be noted that this class is extremely vast and that the examples should not be expected to cover even a substantial fraction of the additional challenges raised by such games.
3.1. One-Player Combinatorial Games
One-player perfect information games are simply combinatorial optimization problems involving the xent function.
assign(s=story())
elicit(t, 10)
ensure("no common words between"+s+"&"+t)
# asks for a nontrivial "summary" of s
reward(xed(s|t))
assign(s1=story(),s2=story(),s3=story())
elicit(t, 10)
ensure("no common words between"+s1+s2+s3+"&"+t))
# asks for a nontrivial "joint summary" of s1, s2, s3
reward(xed(s1|t)+xed(s2|t)+xed(s3|t))
assign(s1=story(), s2=story())
elicit(t, 10)
ensure("no common words between"+t+"&"+s1+s2)
reward(xed(t|s2+s1)+xed(s2|s1+t)+xed(s1|t+s2))
Figure 3.1. Three human-friendly one-player Xent Games (played by $black$, in perfect information mode as per the default)
Figure 3.2. XGD illustrations of the games of Figure 3.1
In Figure 3.1, we present three human-friendly optimization games, which involve the $xed$ function (with $xed\left(s|t\right)=xent\left(s\right)-xent\left(s|t\right)$). Note that the first game is plausibly something that would be a desirable output from an LLM, but that it is actually hard to solve (taking a small temperature leads to a greedy approximation of it). Through substantial effort to try on human players,
assign(s=story())
elicit(t, 10)
assign(s1=s//t,s2=s%t)
# rewards the simplest cut possible
# that makes what comes before
# given what comes after
reward(xed(s1|s2)-xent(t))
assign(s=story())
elicit(t, 10)
assign(s1=s//t,s2=s%t)
# rewards the cut that makes s1
# much more likely to follow s2 than vice versa
reward(xed(s1|s2)-xed(s2|s1))
assign(s1=story(), s2=story())
elicit(t1, 10)
ensure("no common words between"+s1+s2+"&"+t1)
elicit(t2, 10)
ensure("no common words between"+s1+s2+"&"+t2)
# rewards maximally unrelated prefixes
# that are good prefixed for both stories
reward(xed(s1|t1)+xed(s2|t2)+xed(s1|t2)+xed(s2|t1)-xed(t1|t2)-xed(t2|t1))
assign(s=story())
elicit(t, 10)
# find a good prefix t for s
# that is unlikely given s
reward(xed(s|t)-xed(t|s)-xent(t))
assign(s1=story(), s2=story())
elicit(t1, 10)
ensure("no common words between"+t1+"&"+s1+s2)
elicit(t2, 10)
ensure("no common words between"+t2+"&"+s1+s2)
# make the story s1->t1->t2->s2->s1 maximally
# plausible (with the no common words constraint)
reward(xed(s1|t2+s2+t1)+xed(t2|s2+t1+s1))
reward(xed(s2|t1+s1+t2)+xed(t1|s1+t2+s2))
assign(s1=story(), s2=story())
elicit(t1, 10)
elicit(t2, 10)
# find well-explained continuations t1, t2 that are not explaining
# such that t1 is an explanation for t2 but not vice versa
reward(xed(t1|s1)-xed(s1|t1)+xed(t2|s2)-xed(s2|t2))
reward(xed(t2|s1)-xed(s1|t2)+xed(t1|s2)-xed(s2|t1))
reward(xed(t2|t1)-xed(t1|t2))
Figure 3.3. Some simple examples of one-player combinatorial Xent Games
In Figure 3.3, we present a small collection of games each of which comes with slightly different interpretation, despite fairly short codes and limited number of involved variables.
Figure 3.4. XGD representations of the games of Figure 3.3
3.2. Two-Player Zero-Sum Combinatorial Games
3.2.1. Interception/Selection Games
assign(s=story())
elicit(white, t, 20)
elicit(black, t1, 10)
ensure("No common words between" + t + "&" + t1)
# white tries to intercept a priori
# the words that black could use
reward(black, xed(s|t1)+xed(t1|s))
assign(s1=story(), s2=story())
elicit(white, t1, 10)
elicit(white, t2, 10)
elicit(black, t, 10)
# black (who receives the negative of whites score)
# must try to "derail" the story
# that white is trying to build,
# while balancing local consistency
reward(white, xent(s1+t1+t+t2+s2)-xent(s1+t1+t)-xent(t+t2+s2)+xent(t))
Figure 3.5. Interception Games
Figure 3.6. XGD representations of the games of Figure 3.5
Some interesting classes of zero-sum perfect information games are interception games, where one player must anticipate the other player’s moves and hinder their possibilities.
3.2.2. Chess
Another interesting class of games are (variants of) sophisticated games that are (typically) expected to be known to some degree by the judge model $\mathcal{J}$. For instance, the game of chess is in fact very easy to formulate as a Xent Game, given a sophisticated enough judge model able to recognize valid chess moves.
assign(s="")
beacon(flag_1)
elicit(white, t, 4)
ensure("Is valid chess game, black to play or game over:" + s + t)
assign(s=s+t)
elicit(black, t, 4)
ensure("Is valid chess game, white to play or game over:" + s + t)
assign(s=s+t)
replay(flag_1, 20)
reward(white, xent("In the game" + s + "the most likely winner is black"))
reward(black, xent("In the game" + s + "the most likely winner is white"))
assign(s="")
beacon(flag_1)
elicit(white, t, 6)
ensure("Is valid chess game, black to move or game over (specify winner):" + s+t)
assign(s=s+t)
elicit(black, t, 6)
ensure("Is valid chess game, white to move or game over (specify winner):" + s+t)
assign(s=s+t)
replay(flag_1, 5000)
reward(white, xed(s // "#", "white="))
reward(black, xed(s // "#", "black="))
Figure 3.7. Naive Chess and Chess
As described in Figure 3.7, a naive way to play chess using Xent Games is to play 20 moves and then as a judge model to decide who is winning, and use the cross-entropy difference between the estimates as a way; the judge must enforce the validity of moves at every step. Alternatively, the full game of chess can be implemented, but since the theoretical upper bound on chess game lengths is around 5000, the game must run for all these steps.
Figure 3.8. XGD representations of the Naive Chess and Chess games of Figure 3.7
3.2.3. Naive Sprig: Concise Mathematical Proof Debate
Another interesting class of games revolves around (concise) mathematical proofs. The naive idea is to build a proof that can be debated by two parties (inspired by e.g. the Sprig protocol [CGHCL21, CGHCL24]). Note that for this game to correspond in reality to true mathematical debates, a very strong judge model would be needed.
assign(s=story("Get a provable elementary mathematical statement"),s3=story("Get three typical one-line math arguments"))
assign(s0="Here is a proof debate about the statement:" + s)
assign(t0="")
beacon(flag_1)
elicit(white, t, 100)
ensure(xent(t|s0)<xent(s3))
assign(s0=s0+"white proposes to use the following prove the last statement:" + t)
elicit(black, t0, 100)
elicit(black, s, 100)
ensure(xent(s|s0)<xent(s3))
ensure(s + " is a possible point of contention about the debate " + s0)
assign(s0=s0+"black raises a point of contention about " + s0)
replay(flag_1, 100)
elicit(black, t0, 100)
ensure(t0 + "is whites last answer about a point of contention in the debate")
reward(white, xent(t0 + "is a correct elementary mathematical proof w/o missing details: false"))
reward(black, xent(t0 + "is a correct elementary mathematical proof w/o missing details: true"))
Figure 3.9. Mathematical debate
3.3. Imperfect Information Games
Beyond the perfect information setup, the imperfect information games bring considerably more sophistication for the same space of allowed moves. We present a few games that are illustrative of the challenges associated with imperfect information, involving in particular information sharding, coordination, guessing, and bluff.
3.3.1. Secret Sharing
A simple example of an imperfect-information game is secret sharing (Figure 3.10), where Alice must split a secret into Carol’s and David’s share so that neither can find the secret, but so that Bob who has both shares can find it. Notice that this kind of problems have simple cryptographic solutions, but that playing such a game with LLMs can give interesting results.
assign(s=story())
reveal(s, alice)
elicit(alice, s2)
elicit(alice, s3)
reveal(bob, s2)
reveal(carol, s3)
reveal(david, s2)
reveal(david, s3)
elicit(david, t0)
elicit(bob, t2)
elicit(carol, t3)
reward(bob, xed(s|t2))
reward(carol, xed(s|t3))
reward(david, xed(s|t0))
reward(alice, xed(s|t0)+xed(s|t0)-xed(s|t2)-xed(s|t3))
Figure 3.10. Secret Sharing Game
Figure 3.11. XGD representation of the Secret Sharing Game of Figure 3.10
3.3.2. Coordination
Another example of imperfect information game is coordination (Figure 3.12), where Bob and Carol have access to the same information, and must produce the closest possible guesses.
assign(s=story())
reveal(alice, s)
reveal(bob, s)
reveal(carol, s)
elicit(alice, t1, 10)
elicit(bob, t2, 10)
elicit(carol, t3, 10)
reward(alice, xed(s|t1)+xed(s|t1)-xed(t1|x)+xed(t1|t2)+xed(t1|t3)-xent(t1))
reward(bob, xed(s|t2)+xed(s|t2)-xed(t2|x)+xed(t2|t1)+xed(t2|t3)-xent(t2))
reward(carol, xed(s|t3)+xed(s|t3)-xed(t3|x)+xed(t3|t1)+xed(t3|t2)-xent(t3))
Figure 3.12. Coordination Game
3.3.3. Guessing Games
Finally, a very rich family of games are guessing games, where one or several players try to learn information by repeated interactions (see Figure 3.13).
assign(s=story(8))
beacon(flag_1)
elicit(alice, t, 8)
reward(alice, xed(t|s))
replay(flag_1, 100)
assign(s=story(), s0=story(4))
beacon(flag_1)
elicit(alice, t, 4)
reward(alice, xed(s|t))
reward(alice, -xed(s|t))
replay(flag_1, 100)
reward(alice, xed(t|s0)+xed(s0|t))
assign(s=story())
reveal(alice, s)
elicit(alice, t, 10)
reveal(bob, t)
elicit(bob, s0, 10)
reward(bob, xed(s0|t0))
assign(s=story(), s0=story())
reveal(alice, s)
reveal(alice, s1)
elicit(alice, t, 10)
reveal(carol, t)
elicit(carol, t0, 10)
# carol has the power to strip the message from its content
assign(t=t//t0)
reveal(bob, t)
reward(alice, xed(s|t))
elicit(bob, t1, 10)
# alice and bob get double points for passing information about s and for s0
reward(alice, xed(s|t1)+xed(t1|s)+xed(s|t1)+xed(t1|s)+xed(s0|t1)+xed(t1|s0))
reward(bob, xed(s|t1)+xed(t1|s)+xed(s|t1)+xed(t1|s)+xed(s0|t1)+xed(t1|s0))
# carol is incentivized to let information about s pass
# she is doubly punished for information about s1
reward(carol, xed(s|t1)+xed(t1|s)-xed(s1|t1)-xed(t1|s1)-xed(s1|t1)-xed(t1|s1))
Figure 3.13. Guessing Games
4. Xent Game Space: Model Play Properties
Being made of text interactions mediated by LLM judges, Xent Games introduced in Section 2 are naturally designed to be played by LLM-based agents. In this section, we focus on the relationship between Xent Game play and LLMs.
4.1. Xent Game Eval Mode
By definition a Xent Game $G$, beyond its game operation code, involves some metadata $D$ including ‘linking’ metadata:
  • the specification of the judge model $\mathcal{J}$;
  • the NPC models linked to the set $\mathcal{A}\left(G\right)$ of active players (i.e. the ones which are involved in at least one $elicit$ statement).
When using a Xent Game to evaluate a model $\mathcal{M}$, the idea is to ‘freeze’ a game $\left(G,D\right)$, pick an active player $P\in\mathcal{A}\left(G\right)$ in the game $G$ and to consider the family of games $\left(G,D|_{P=\mathcal{M}}\right)$ where $D|_{P=\mathcal{M}}$ denotes the linking of $P$ with the model $\mathcal{M}$, overriding the linking of $\mathcal{P}$ in $D$, i.e. forcing the player $P$ to become the main character played by $\mathcal{M}$. Informally, this corresponds to the following simple idea: consider a game $G$, pick a player $P$ and measure how the model plays the role of the ‘character’ $P$.
Typically, running a Xent Game in Eval Mode furthermore involves specifying a random ‘map-seed’ $\sigma$ used to set the instance random data (unless we deal with a game with deterministic instances, e.g. like Chess or Go), and in the outputs of the players except the model being evaluated; we typically use a different ‘play-seed’ $\varsigma$.
Hence, when discussing the (unnormalized) score of a given model $\mathcal{M}$ on a game $G$, we consider the averaging over random values $\sigma$ of the reward in the game $\left(G,D|_{P=\mathcal{M},\mathcal{S}}\right)$ where $P$ is overridden to be played by $\mathcal{M}$ and the seed is set to random values $\sigma_{i}$ for $i=1,\ldots,N$:
(4.1) $S_{G}\left(\mathcal{M}\right):=\frac{1}{N}\sum_{i=1}^{N}S_{G}\left(G,\sigma_{i},\varsigma_{i}\right).$
Remark 27.
This definition is closely related to that of the $arms$ definition given in Section 4.2 below. It is important in practice, for us to not let $N\to\infty$ (which would yield an expectation), as this leads to undesirable features (see Remark 34).
Remark 28.
In order to allow for unified normalization across various games, game maps and game settings, the game scores should be normalized using a base model (see Section 5.2 below).
4.1.1. Game Representation and Variance
When averaging over map seeds in practice, it is important to take into account the variance of rewards with respect to the map seed; this sometimes singles out certain formulations of games that otherwise have the same optimal solutions and strategies.
For instance, the basic reverse prompting game (Example 14 above) admits the following equivalent representations: when the reward $-xent\left(s|t\right)$ is given to $white$ and when $xed\left(s|t\right)=-xent\left(s|t\right)+xent\left(s\right)$ is given to $white$: the scores as a function of the move $t$ just differ by a constant. The latter formulation is typically to be preferred, as its variance with respect to the randomness of $s$ (induced by the map-seed $\sigma$) is lower; the reliance of the game designs presented in Section 3 upon the $xed$ function is largely motivated by this consideration.
4.1.2. Well-Posedness
A notion that is related to the one discussed in Section 4.1.1 above is that of well-posedness. This can be viewed as an analog to the notion of boundedness in optimization: trying to maximize a function that is not bounded from above cannot yield meaningful results, even if in practice values may be bounded by machine precision (whatever result we will get will reflect something about the machine running the computations rather than about the problem). In the setting of Xent Games, rewards on any game instance are naturally bounded by the bounded sizes of the strings, but that does not imply solutions are anyhow meaningful, as they may be more reflective of irrelevant model artifacts rather than anything else.
Remark 29.
A good example of this is the simple game that consists in maximizing $xent\left(s\right)$ over all strings $s$ of a certain length: this leads the model to ‘tap’ into words with extremely low probability whose estimates have usually no reason to have been precisely calibrated, and the results will typically look like random text.
Precisely defining the notion of well-posedness for Xent Games is in fact not trivial: as much as it is easy to come up with examples of ill-posed games, such as the one of Remark 29, distinguishing between well-posed and ill-posed games can be difficult. A simple intuitive criterion is that gameplay should not be affected by very small probabilities, i.e. that if we put a little bit of noise in the probability vector, the game-play should not be substantially altered (either by the modification of allowed moves or by the scoring of the outcome). While this could be formulated into a precise definition, the following is essentially equivalent and much simpler in practice:
Definition 30.
For $\epsilon>0$, we say a Xent Game $G$ (in eval mode) is $\epsilon$-well-posed if the main player’s optimal strategy is not affected by $\epsilon$-clipping of probabilities, i.e. replacing $x\mapsto-\log x$ by $x\mapsto-\max\left(\log x,\log\epsilon\right)$ in the cross-entropy.
Remark 31.
In practice, a reasonable (though fairly arbitrary) value we will take for $\epsilon$ is $1/|\mathcal{V}|$, i.e. we will disregard games where optimal play relies on tokens that are less likely than what a random uniform guess would give.
Remark 32.
By this definition, it is not hard to see that the game of Remark 29 is ill-posed, as is the paradoxical game of Example 16 (when in eval mode with $white$ or $black$).
4.1.3. Game Repetition
An important feature of learning to play games is repetition, i.e. learning from mistakes, and more generally understanding the interplay between a game play history and the rules of the games to optimize one’s strategy, by putting all of this information as a context for the model to play. A useful setting in practice is to allow the to replay of games by the main character played by the model $\mathcal{M}$, while keeping the same map-seed and the same play-seed over the various replays, giving as context the past games to $\mathcal{M}$, while keeping no past game context to the other NPC agents. This setup, which advantages $\mathcal{M}$ over the NPC players, can be expected to give an increased score as the attempts are repeated for a reasonable class of games: as discussed in Section 4.2.1 below, we call these the few-shot playable games.
4.1.4. Game Fine-Tuning
Another important way in which we can see a model getting better at games is by processing examples and adjusting its parameters on them, i.e. to perform some fine-tuning based on reinforcement learning principles. We will abstract the many subtle details involved with such a process by merely assuming that we are given, for some model $\mathcal{M}$, a process $\mathcal{F}$ that, given a game $G$ (with an assignment of $\mathcal{M}$ to some player $P$ in $G$), outputs a model $\mathcal{F}_{G}^{\mathcal{T}}\mathcal{M}$ corresponding to ‘$\mathcal{M}$ fine-tuned to play better at $G$’ for some time $\mathcal{T}$.
While we intentionally avoid setting specific details about the process, it is useful to take as an example the one developed in Section 4.3.2. Note that we do not require that the model has actually gotten better at $G$: as will be discussed in Section 4.2 below, this corresponds to the family of fine-tuning playable games.
It should be noted that a fine-tuning mechanism should not be assumed to exist for all models $\mathcal{M}$: for instance, some closed models can be just available in inference mode and not be fine-tunable. In this paper, we simply required the existence of a sufficiently strong base model $\mathcal{M}_{\mathcal{B}}$ for which a fine-tuning mechanism is available, that makes enough games playable (see Section 6 below).
4.2. Playability: Few-Shot and Fine-Tuning Definitions
Central to our investigation of well-posed Xent Games is the notion of playability, which informally corresponds to a quality of certain games that models can get improved performance on, either by repeated play context (few-shot definition) or by reinforcement learning (fine-tuning definition). This should be thought of in opposition to non-playable games, which contain for instance as a non-trivial example the breaking of a secure code (even with a small secret length), as one either succeeds or fails and there is no intermediate reward and no sense that we get closer to succeeding.
4.2.1. Few-Shot Definition
A remarkable ability of modern LLMs is their so-called few-shot learning abilities [OAI2020, KGRMI22]: from a limited number of feedbacks directly put in their context window, models are often able to substantially improve their responses. The notion of few-shot playability by a model $\mathcal{M}$ refers to those games where the few-shot abilities are reflected in game-play:
Definition 33.
For a game $G$ and $N$ random samples (of map seeds and play seeds), we define the average running max score $\mathrm{arms}_{N,k}\left(G,\mathcal{M}\right)$ as the (random) sequence defined by
$\mathrm{arms}_{N,k}\left(G,\mathcal{M}\right)=\frac{1}{N}\sum_{\ell=1}^{N}\max_{j=1,\ldots,k}\left(S_{j}^{\left(\ell\right)}\right),$
where $S_{j}^{\left(\ell\right)}$ is the score at the $j$-th iteration over the sample $\ell$.
A well-posed game $G$ is $\left(N,\epsilon,m\right)$-playable in few-shot mode by a model $\mathcal{M}$, if with probability $1$-$\epsilon$, the score $\mathrm{arms}_{N,k}\left(G,\mathcal{M}\right)$ is strictly increasing for at least $m$ steps. We say that a playable game is $\left(N,\epsilon,\theta,m\right)$ playable if there exists $m$ thresholds $k_{1}<\cdots<k_{m}$ such that $\mathrm{arms}_{N,k}\left(G,\mathcal{M}\right)$ exceeds $b+j\theta$ for the first time at $k_{j}$ for $j=1,\ldots,m$ where $b$ is the expected score at the first iteration.
In practice, we will assume that some values of $N,\epsilon,m$ are set. In the numerical experiments below (Section 4.2.2), we take $N=20$ and $m=30$.
Remark 34.
This definition means that in practice, averaging over a sufficiently large number of samples, we do see the running max score increase substantially for a while (i.e. at every step, the model has done better in one sample). We choose this definition rather than that of an expectation as the latter is sensitive to very unlikely (e.g. breaking a secure cipher is not a playable game, yet the probability of cracking it is technically nonzero at every step).
As we will see, in Section 5, few-shot playability is an important a priori property for benchmarking: the benchmarking score of a game will be defined relative to the times when a base model $\mathcal{M}_{\mathcal{B}}$ crosses $\theta$-levels for a game’s repeated play.
4.2.2. Numerical Example: Benchmarking with Single-Text
As an example, the single-text game presented in Figure 3.1 is a good example of a playable game that can be used to compare various model performances. Our results confirm that current state-of-the-art LLMs are indeed capable of playing a Xent Game written in XGL and playable in the sense of Definition 33: they tend to improve over iterations.
The initial benchmarking results have a number of interesting features. That being said, the results clearly demonstrate that current models are able to improve their performance on a well-posed game when given the chance to provide new answers after seeing the results of their previous choices.
The following plots show performance of GPT-4.1, Claude Opus 4, Gemini 2.5 Pro, and Grok 3 when playing the single-text game posed in the Xent Game Language in an iterated scenario using GPT2 as the judge model. Each test case is the single text game with a different generated string. While these specific test cases show clearly improvement across the set of test models, often results would be more mixed, showing flat or even occasionally negative performance given more information.
Figure 4.1. Performances for four flagship models playing single text Xent Game (with $N=20$, running for $32$ iterations): the top plot shows the averaged scores (see Expression 4.1) and the bottom plot shows the $arms$ (see Definition 33). We see that the $arms$ keeps improving for all models up to $k=17$.
4.2.3. Fine-Tuning Definition
The other, somehow deeper and more general, notion of playability pertains to the ability of models to get better by fine-tuning performed for long enough. It should be noted that closed models are not always available in fine-tuning mode, and so that this notion is strictly less applicable than the few-shot playability one:
Definition 35.
A well-posed game $G$ is playable in fine-tuning mode by a model $\mathcal{M}$ if for some time $\mathcal{T}>0$, the model $\mathcal{F}_{G}^{\mathcal{T}}\left(\mathcal{M}\right)$ is strictly better in expectation than the model $\mathcal{M}$ at $G$.
Remark 36.
It is generally expected that if a game is playable in fine-tuning mode, it is also playable in zero-shot mode.
As we will see in Section 4.3, the fine-tuning playability is crucial for us to relate games with one another via the so-called transfer value. This notion of transfer value then allows us to build a geometric structure on the space of Xent Games, which we rely upon to construct relevant Xent Measures.
4.3. Transfer Value
As discussed in Section 4.2, a playable game is one where a player’s strategy can be adjusted by experience, leading to a higher score; in other words, the model $\mathcal{M}$ is learning something. But is it worth playing the game, i.e. is the learned information valuable? Surely, this is in large part very subjective. In this subsection, we will define the value in terms of the increased abilities of an agent fine-tuned to play that game. This of course shifts the subjectivity further: what is the end-goal in terms of game playing utility? We defer the investigation of the latter problem to Section 5.
4.3.1. Transfer Value of a Game
From the fine-tuning mechanism defined in Section 4.2.3, we can now ask the question: how much does playing a game for a bit improve an agent’s skill on another game? This leads to define the value of a playable game:
Definition 37.
Given a model $\mathcal{M}$ and a fine-tuning mechanism $\mathcal{F}$ and a fine-tuning time $\mathcal{T}>0$ (i.e. a token processing budget), we define the transfer value (relative to $\mathcal{F)}$ of a playable Xent Game $G_{1}$ for another playable game $G_{2}$ by
$\mathcal{V}_{G_{1}}\left(G_{2}\right)=\frac{S_{G_{2}}\left(\mathcal{F}_{G_{1}}^{\mathcal{T}}\left[\mathcal{M}\right]\right)-S_{G_{2}}\left(\mathcal{M}\right)}{S_{G_{1}}\left(\mathcal{F}_{G_{1}}^{\mathcal{T}}\left[\mathcal{M}\right]\right)-S_{G_{1}}\left(\mathcal{M}\right)},$
where $S_{G}$ denotes the expected score of a model when playing $G$ (averaged over seeds).
Remark 38.
The transfer value should simply be understood as ‘how much value does it bring towards playing $G_{2}$ to play $G_{1}$?’ (normalized by for how long one plays $G_{1}$). Note that there is no reason to expect symmetry, i.e. $\mathcal{V}_{G_{1}}\left(G_{2}\right)$ is generally not equal to $\mathcal{V}_{G_{2}}\left(G_{1}\right)$.
Remark 39.
As much as we will often take for granted that $\mathcal{M},\mathcal{F}$, and $\mathcal{T}$ have been suitably set, working with transfer values, it is important to remind ourselves that $\mathcal{V}$ depends a priori on them, and that a ‘good setting’ for $\mathcal{M},\mathcal{F},\mathcal{T}$ has been found. It is however important to keep in mind that we can still expect that the final outputs of our constructions (see Sections 5 and 6 below) to not depend on $\mathcal{M},\mathcal{F},\mathcal{T}$: we expect all ‘good’ $\mathcal{M},\mathcal{F},\mathcal{T}$ settings give roughly the same general ability measures (see Section 6.2.3 for a discussion of this universality idea).
Remark 40.
The idea of transfer learning across games appears in a number of places in the literature, see e.g. [BaSt07, PBS16, CKATT18]; generally speaking it is reasonable to expect similar games to be related. It is also interesting to note that nonzero transfer value between an imperfect information game and its perfect information counterpart (i.e. where all information is revealed) can be seen experimentally (see [BCK23]).
While the transfer value of a game of course depends on the many details abstracted in the $\mathcal{F}$ mechanism (and of the game metadata, in particular the models linked to the judge and players), it is an interesting question to ask how much the constructions we build from it depend on such details (see Section 6).
4.3.2. Transfer Value: Numerical Example
The transfer value concepts introduced above can be illustrated on a simple triplet of games (see also Section 3.1 for similar games).
# Game Name: Single Text
assign(s=story())
elicit(t, 10)
ensure("no common words between"+s+"&"+t)
# asks for a nontrivial "summary" of s
reward(xed(s|t))
# Game Name: Multi Texts
assign(s1=story(), s2=story(), s3=story())
elicit(t, 10)
ensure("no common words between"+s1+s2+s3+"&"+t))
# asks for a nontrivial "joint summary" of s1, s2, s3
reward(xed(s1|t)+xed(s2|t)+xed(s3|t))
# Game Name: Dex Texts
assign(s1=story(), s2=story())
elicit(t, 10)
ensure("no common words between"+t+"&"+s1+s2)
# asks for a summary of s1, but anti-summary of s2
reward(xed(s1|t) + xed(s1|t) - xed(s2|t))
Figure 4.2. Triplet of one-player Xent Games (played by $black$, in perfect information mode as per default)
These games, while clearly related, have distinct rules and strategies. In particular, the third game utilizes a negative $xed$, meaning that the playing agent is incentivized to find a prefix $t$ which makes $s2$ unlikely.
Experimentally, we show the presence of transfer learning (i.e. nontrivial transfer value) by making three independent fine-tuning (one on each of the three games) of a model (Qwen3 8B in FP8 in our experiment with GPT2-XL as the judge model): for each of the three models being fine-tuned, we evaluate their performance on all three games. The fine-tuning procedure $\mathcal{F}$ is performed using the REINFORCE algorithm [Wil92] on rewards associated with the choices of $\mathcal{M}$ at every $elicit$ call, run with some the AdamW optimizer and learning rate $10^{-6}$ with a token processing budget $\mathcal{T}$, with a corpus of 20k ‘stories’; the evaluation is performed on a corpus of 600 ‘stories’. Our experiments demonstrate a clearly non-trivial transfer value across games.
Figure 4.3. Evaluation results during reinforcement learning for “Single Text”, “Multi Texts”, and “Dex Texts” games, respectively.
These results demonstrate that even in very simple configurations there is clear transfer learning between games occurring and that the transfer rates between games are varied and not symmetrical. We see in particular that for the “Dex Texts” game, we have an asymmetry between transfer values (see Remark 38): for instance, learning “Single Text” helps more for playing “Dex Texts” than learning “Dex Texts” helps for playing “Single Text”.
5. Xent Game Measures: Scopes and Covers
In this section, we construct, Xent Measures, which yield theoretically-motivated LLM benchmarks from Xent Game scopes, which are families of playable Xent Games representing a desirable set of abilities. This will pave the way towards Section 6, which defines general capability scopes, leading to general ability measures.
5.1. Benchmark Goals
In this subsection, we first review some works on LLM benchmarking, in particular towards measuring general abilities.
5.1.1. From specific to general abilities: benchmarking
Traditionally, the benchmarking of LLMs tests their explicit knowledge using a set of questions (that are typically made by humans) with e.g. multiple-choice answers. Designing such benchmarks is of course very contingent on specific areas; there is also naturally the risk of data contamination. See e.g. [LBL23, PCDSC24, CWWWX23] for surveys on the works.
Interesting examples of benchmarks that attempt to go towards more general abilities include for instance [Cho19, CKKLP25]. A very promising line of work that generally attempts at evaluating LLM performance is via the creation of environments to evaluate them, see e.g. [XDCGZ24, SSSJ25, CKATT18], though a principled way to extract a benchmark from such environment is not yet available. The idea to use games has been popularized in e.g. [CZJGS24].
In spite of there being a plethora of mechanisms to benchmark LLMs, the interpretations and theoretical underpinnings of such benchmarks are arguably rarely clear. In fact, one could even say that it is not very clear what one expects from a benchmark, besides answering the informal (and very underspecified) question of which LLMs are ‘better, more capable’. Since these notions are not very well defined a priori, if one asks the question ‘what is expected from a new benchmark’, a key element one expects to inform us along some dimension that is not captured accurately by other benchmarks, but that one will recognize (perhaps a posteriori) as being driven by useful skills.
While the practical successes yielded in part by the current state of the art is undeniable (given that, ultimately, most are able to recognize spectacular abilities of LLMs, without usually being able to define them), this state of affairs makes it naturally hard to ascribe a specific meaning to a benchmark score or to clearly identify needs for future benchmark scores.
5.1.2. Vision and Goals
The vision that we provide is in fact very simple: we would like benchmark results to have simple interpretations in terms of concrete abilities on (hopefully quite general) families of tasks, i.e. to provide us with some statistics telling us ‘which fraction of tasks within a certain scope can be accomplished at what skill level’. Such measures can thus be used to compare capabilities of models with one another, and also steer the development of LLM abilities to fill some gaps in capabilities. In some sense, if we can find a fair sample of tasks that are roughly equivalent in terms of relevance, the output benchmark should be a histogram of the scores over the various tasks: should the need to further summarize the model’s abilities in one or several scalars arise, means, medians, and more general moments and quantiles can give a good idea of the situation.
Remark 41.
The idea to use text-based games as means to evaluate LLM abilities has been considered in [TKM23].
5.2. Score Normalization
As discussed above, the goal of benchmarking is naturally to compare various models (whether of different architectures, datasets or simply fine-tuning of a given model). Naturally, each playable Xent Game $G$ instance can be used to measure various models; given the diversity of such games, the question of score normalization, in particular across games, roles, and seeds, is not obvious.
5.2.1. Base Model and Few-Shot Mode
Related to score normalization is that of units: what are the units in which we aim to understand the performance of models playing Xent Games? While the Xent Games have the natural advantage of having scores expressed in bits, surely the bits are not each of the same value; even within the family of games that are playable for some model, some are much harder to play than others.
For instance, going from 0.95 win rate at playing chess against an agent to 0.999 win rate could be interpreted as a very drastic increase in ability, while it would correspond to a few percent of score increase (much more than going from a rate of 0.5 to 0.55, say).
In order to normalize scores, a simple solution is to identify a sufficiently strong base model $\mathcal{M}_{\mathcal{B}}$, find reasonable values of $N$ (number of samples), $\epsilon>0$ (tolerated ‘failure probability’), $\theta$ (increment value) and $m\geqslant 2$ (number of increments), and to consider the set of $\left(N,\epsilon,\theta,m\right)$-playable games relative to $\mathcal{M}_{\mathcal{B}}$ (see Definition 4.2.1), which we will simply call (unless specified otherwise) the playable Xent Games.
From there, the performance of any model $\mathcal{M}$ can be measured in terms of how fast $\mathcal{M}$ (in terms of iterations) can reach the level of $\mathcal{M}_{\mathcal{B}}$, and hence yields results that can be measured in terms of work done per iteration compared to $\mathcal{M}_{\mathcal{B}}$.
This measure of a model’s abilities on a game in economic terms is particularly useful to capture its performance on a task. If we have a good estimate of a model’s score, we will hence have a good estimate on its progression curve when playing a new game, and not only on its eventual performance.
5.2.2. Normalized Score Definition
As suggested in Section 5.2.1, we leverage the concept of playability to obtain a unified normalization scheme for Xent Games, once a base model $\mathcal{M}_{\mathcal{B}}$ is picked.
Definition 42.
Let $G$ be $\left(N,\epsilon,\theta,m\right)$-playable by a base model $\mathcal{M}_{\mathcal{B}}$ for some fixed values of $\left(N,\epsilon,\theta,m\right)$ (see Section 4.2.1 above), with base mean score $b$ at the first iteration times $k_{j}$ when $j$ exceeds $b+j\alpha$ for $j=1,\ldots,m$. Then for any $\ell\leqslant k_{m}$ and any model $\mathcal{M}$ yielding $arms_{N,k}$ scores (see Section 4.2.1 above) at the various iterations, the normalized score of $\mathcal{M}$, denoted $narms_{N}$, at time $\tau=\sup\left\{k\in\left\{1,\ldots,\ell\right\}:arms_{N,k}\leq b+m\alpha\right\}$ is defined by:
  • If $\tau=\ell$ (i.e. $\mathcal{M}$ has not outpaced $\mathcal{M}_{\mathcal{B}}$ in $\ell$ steps) and $arms_{N,\tau}=b+j\alpha$ for some $j\leq m$, then $narms_{N}$ is exactly $j$; anything in between is the unique affine interpolation that is continuous in with respect to $arms_{N,\tau}$.
  • If $1<\tau<\ell$ (i.e. $\mathcal{M}$ has outpaced $\mathcal{M}_{\mathcal{B}}$ in $\ell$ steps) and $arms_{N,\tau}$ is exactly $b+m\alpha$, the $arms_{N}$ score is $m\frac{\ell}{\tau}$; anything in between is given by the unique affine interpolation that is continuous in $arms_{N,\tau}$.
Informally, the normalized score consists of the number of play steps of $\mathcal{M}_{\mathcal{B}}$ that $\mathcal{M}$ is able to catch up with per step.
5.2.3. Multi-Game Benchmark Design
In Section 5.2, we presented the question that a unified normalization aims to answer the following: What is the performance of a model playing a game, relative to a base model?
Following this route, we can now discuss the question of a multi-game benchmark normalization. Again, a fairly simple question can be asked: if a model is supposed to play a large number of games, if we run it for a certain number of steps on each game, what does it deliver on average, per game, relative to a base model $\mathcal{M}_{\mathcal{B}}$? The idea of using an averaging measure again follows from economic considerations: running costs are naturally additive. In addition to the above motivation, averages (or perhaps more generally weighted averages) have the advantage of being maximally sensitive to individual game benchmarks (if we use weighted averages, we can explicitly say how important each game is).
The problem of selection of the games appears however to be substantially more subjective and challenging: we discuss this question in Section 5, and an answer based on an evolution-inspired meta-game will be provided in Section 6.
5.3. Benchmarks from Xent Measures
As a way to propose a solution to the benchmarking problem raised in Section 5.1 above, from to the structure of the Xent Game space 2.3.3 it is tempting to suggest to rely on benchmarks of the following forms for a suitable (finite) family of games $\Gamma$:
Definition 43.
For a family of games $\Gamma$ and a normalized score function $S^{*}$ (such as the $narms$ defined in Section 5.2.1), the Xent (probability) measure $\mu_{\Gamma}\left(\mathcal{M}\right)$, is defined as $\frac{1}{\left|\Gamma\right|}\sum_{G\in\Gamma}\delta_{S^{*}\left(\mathcal{M}\right)},$ where $\delta_{x}$ denotes the Dirac mass at $x\in\mathbb{R}$.
Remark 44.
The above should simply be understood as the histogram of the set of values $S^{*}\left(\mathcal{M}\right)$ over $G\in\Gamma$ (accounting for repeated values).
In order for $\mu_{\Gamma}$ to serve as the basis of a ‘good’ benchmark for a given scope, the games in $\Gamma$ must hence be chosen in such a way that:
  • The family is large enough to satisfactorily cover the diversity of the scope.
  • The covering is ‘fair’, i.e. there is no ‘over-representation’ of some games on the scope.
In Section 6.2 below, we will see that the notion of transfer value defined above (Section 4.3) can be used to achieve such goals.
5.4. Xent Measure from Scope
The idea behind the definition of Xent Measures from scopes is simple, and very related to meta-learning (see e.g. [Bax00] for foundations, [SYSKZ24] for a recent survey about LLMs, and [BVLFW25] for a survey about meta-reinforcement learning): given a scope $\Sigma$, find a minimal covering family $\Gamma$ such that any game in $G\in\Sigma$ is ‘close enough’ to a game in $G^{\prime}\in\Gamma$ (in the sense that $\mathcal{V}_{G^{\prime}}\left(G\right)$ is large enough): in other words, we want a minimal set of games $\Gamma$ such that if we had to play any game $G\in\Sigma$, and if we are able to play the games in $\Gamma$, we could quickly fine-tune our model to play $G$.
A minimal covering family $\Gamma$ naturally yields a capability benchmark for LLMs on the scope $\Sigma$: for a model $\mathcal{M}$, consider as the (histogram) of the scores $S_{G}\left(\mathcal{M}\right)$ for $G\in\Gamma$. Intuitively, the idea is that if $\Sigma$ is dense, each $G\in\Gamma$ plays a similar role, and can thus be considered interchangeably, with the minimality guaranteeing ‘no over-representation’ of certain regions of the game space: the histogram of the scores then gives us a good idea of how well $\mathcal{M}$ can be expected to perform on a random game sampled ‘near $\Sigma$’ (i.e. in a region well covered by $\Sigma$).
Despite the fact that the following definition is quite theoretical (even for very simple cases, computing the optimal cover is NP-hard), it is reasonable to aim for approximations (and find good covers).
Definition 45.
Consider a scope $\Sigma$, and a space of games $\Upsilon$.
  • For a family of Xent Games $\Gamma\subset\Upsilon$, we denote by $\Delta_{\Gamma}\left(\Sigma\right)$ the transfer cover of $\Sigma$, defined as $\min_{G\in\Sigma}\max_{G^{\prime}\in\Gamma}\mathcal{V}_{G^{\prime}}\left(G\right)$ : this represents how much $\Sigma$ is guaranteed to be covered by $\Gamma$.
  • We call a family of Xent Games $\Gamma$ an $\eta$-cover if $\Delta_{\Gamma}\left(\Sigma\right)\geq\eta$.
  • We say that an $\eta$-cover of $\Sigma$ is optimal if its cardinality is minimal and if its transfer cover is maximal for its cardinality, and we denote such a cover $\Gamma_{\Sigma}^{\eta}$.
Remark 46.
We have that $\Gamma_{\Sigma}^{\eta}$ is in some sense a ‘skeleton’ of $\Sigma$, i.e. a version of $\Sigma$ stripped of duplicate games, and possibly games replaced by ‘leaner versions of themselves, for games $G$ such that there exists a $G^{\prime}$ with $\mathcal{V}_{G^{\prime}}\left(G\right)>\mathcal{V}_{G}\left(G\right)$.
From the above, we obtain:
Definition 47.
For a scope $\Sigma$, and a cover $\eta$ we call benchmark measure associated with $\left(\Sigma,\eta\right)$ the Xent Measure $\mu_{\Sigma}^{\eta}:=\mu_{\Gamma_{\Sigma}^{\eta}}$.
Remark 48.
Even if the scope $\Sigma$ is large, most of the interesting covers should correspond to relatively small values $\eta$: these correspond to fairly coarse covers.
5.5. Scope Selection
From Section 5.4, given a finite scope $\Sigma$, a Xent benchmark measure $\mu_{\Sigma}^{\eta}$ can be defined for any $\eta\leq 1$. This leaves open the question of choosing the scope itself. Depending on one’s goals, different types of considerations can apply:
  1. (1)
    We have a set of specific goals that are covered by a set of games, such as the ones introduced in Section 3.
  2. (2)
    We want to highlight differences in capabilities between various models: we would like to select games that are maximally discriminative between models.
  3. (3)
    We are aiming to measure general abilities, in particular how a model will perform in a setting that is not known a priori.
These considerations can apply separately or simultaneously: for instance, we may have a set of pre-defined goals, but would like to expand in generality around those, or may want to set goals to allow a model to differentiate itself from other models.
Since the key problem raised by the Items 1 and 2 is to construct a specific list of games (whether hand-crafted or picked according to a feature selection process), we see that in any case the problem is the following: to answer general abilities question 3, starting from a list of specific games (see Problem 49) below.
5.5.1. Towards a General Ability Measure
A naive approach would be to say that since ‘general abilities’ are about all the games, we should want to put all games on equal footing; an obvious problem is the infinite size of the resulting scope, making it unlikely we can cover the space with a finite (let alone small) set of games. Intuitively, this makes sense, as it is unlikely that ‘general abilities’ could ever be well summarized by a fixed number of games: for any finite game set, it is very plausible we can find a game that is substantially different from all games in that set.
Giving up on the idea of finite set of games appears to be a promising direction, bringing us closer to something like a curriculum. Naturally, this forgoes the idea of a uniform benchmark measure (see Remark 50 below), and leads us to focus on an order for games that are to make up a benchmark.
Problem 49.
Starting from an initial scope $\Sigma_{0}$, what is a principled way to define a sequence of Xent Measures $\mu_{k}$ which form a good estimate of a model’s general abilities around $\Sigma_{0}$?
Remark 50.
While ordering measures (and scopes) introduces a subjective bias, this seems absolutely necessary for both theoretical and practical reasons: this is similar to the fact that it is impossible to sample an (unbounded) integer uniformly at random.
Remark 51.
In practice, if, for a given model $\mathcal{M}$, the Xent Measures $\mu_{k}\left(\mathcal{M}\right)$ stabilize around some value $\mu_{*}$, it seems reasonable to ascribe that value $\mu_{*}$ as a measure of $\mathcal{M}$ as a ‘general ability measure of $\mathcal{M}$’ (even though the latter may not be definable generally: this is the same kind of criterion used to say that ‘50% of the integers are even’, in spite of the fact that there is no uniform measure on integers).
While this question is naturally very under-specified, we will see in Section 6 that evolution-inspired ideas can lead us towards answers to the Problem 49.
6. Evolution in Game Space
In Section 5, we introduced the idea of benchmarks based on scopes, i.e. families of Xent Games. In Section 5.5.1, we discussed the challenges associated with infinite scores, and formulated Problem 49: how to find, starting from an initial family of games $\Sigma_{0}$ a sequence of measures that can be viewed (thinking of the $k\to\infty$ limit) as an approximation of general abilities (in the sense of Remark 51)?
In this section, we consider ways to explore the space of Xent Games by leveraging its geometric structure. We then propose, from game-theoretic considerations and evolution-inspired ideas, mechanisms to define sequences of measures $\mu_{k}$ that aim at measuring general model capabilities.
6.1. Xent Game Space Geometry
One of the key insights about the Xent Game space is that the games are related via their structures to one another: this is essentially a consequence of the proof of Theorem 25 above. This suggests (informally) that if one has a strong enough base model $\mathcal{M}_{\mathcal{B}}$, one should be able to transform any (reasonable) Xent Game into any other Xent Game by a sequence of moves such that we can transfer some non-trivial amount of skill from game to game (locally, not necessarily across the whole path).
Definition 52.
For $\alpha>0$ and a base model $\mathcal{M}_{\mathcal{B}}$, we write $G_{1}\rightmoon_{\alpha}G_{2}$ if $\mathcal{V}_{G_{1}}\left(G_{2}\right)\geqslant\alpha$ and $G_{1}\leftmoon_{\alpha}G_{2}$ if $\mathcal{V}_{G_{2}}\left(G_{1}\right)\geqslant\alpha$, and call a finite sequence of games $G_{1},\ldots,G_{n}$ an $\alpha$-downstream path if $G_{1}\rightmoon_{a}\cdots\rightmoon_{\alpha}G_{n}$ and an $\alpha$-upstream path if $G_{1}\leftmoon_{\alpha}\cdots\leftmoon_{\alpha}G_{n}$.
Remark 53.
As suggested above, there is no reason to expect a priori $\rightmoon_{\alpha}$ to be a transitive notion.
Definition 54.
For $\alpha>0$ and a collection of games $\Sigma_{0}$, we denote by $\left(\Sigma_{0}\right)_{\alpha}$ the $\alpha$-downstream component defined as the set of games $G$ reached by an $\alpha$-downstream path that starts from a game in $\Sigma_{0}$; similarly, we denote by $\left(\Sigma_{0}\right)^{\alpha}$ the $\alpha$-upstream component defined as the set of games reached by an $\alpha$-upstream starting from a game in $\Sigma_{0}$.
We say that a space of Xent Games is $\alpha$-connected if there is a finite $\Sigma_{0}$ such that the downstream component $\left(\Sigma_{0}\right)_{\alpha}$ is the whole space. In other words, the set of Xent Games is $\alpha$-downstream connected to a finite set of games.
For a ‘sophisticated enough model’, it is reasonable (by the construction of Theorem 25 above) to postulate that all the (playable) Xent Games are related, in the following sense: one can always modify a playable game into another one by a chain that is $\alpha$-downstream and by a chain that is $\alpha$-upstream, for some small enough $\alpha$. This is the content of the following:
Assumption 55.
There exists some $\alpha>0$ such that the space of playable Xent Games is $\alpha$-connected, i.e. that we can reach any playable game from a finite set of basic playable games (e.g. the basic Xent Game) by $\alpha$-downstream paths and also by $\alpha$-upstream paths.
Remark 56.
This may suggest to take $\alpha$ small enough; or we could also instead decide to focus on an $\alpha$-connected component for some a priori fixed $\Sigma_{0}$ (if we don’t care about general abilities, but simply general abilities around some scope).
As the space of Xent Games is formally infinite, the situation may seem analogous to the situation where we need to explore an infinite-dimensional vector space, and it may as a result seem hopeless to ‘exhaust’ it a priori (i.e. without additional structures) with a growing sequence of ‘uniform’ measures on finite sets (as follows from e.g. Baire’s Theorem). As explained in Section 6.1.1 below, we will essentially give up on this analytic point of view to focus on heuristics coming from samples from exploration processes that leverage some weak form of ergodicity: if the measures we find through different runs are consistent for a given model, we postulate that we have a ‘good’ measure of the model’s ability (and we postulate that inconsistencies are unlikely, based on universality assumptions, see Section 6.2.3).
Remark 57.
Once a base model $\mathcal{M}_{\mathcal{B}}$ is fixed, the geometric picture outlined above is most of what we use about playable Xent Games in this section (i.e. we abstract the specific details of Xent Game implementation).
6.1.1. Web Exploration Intuition
If we assume that space of Xent Games is $\alpha$-connected for some $\alpha>0$, we can think of it as an (infinite) graph, we can put a downstream link from $G_{1}$ to $G_{2}$ if $G_{1}\rightmoon_{\alpha}G_{2}$ and an upstream link if $G_{1}\leftmoon_{\alpha}G_{2}$. We can hence leverage this structure to think of the space of games as a kind of ‘inter-game web’ and use this vision as a guiding principle to explore it like a (very) complex network.
Note that unlike the internet, it is a priori very non-trivial to list all the links from a given game; however, we can think that in practice, there is a from a finite list of games $\Sigma$, it is definitely possible to determine which ones are $\alpha$-linked and which ones are not, thus equipping $\Sigma$ with that structure.
Remark 58.
In practice, the way the exploration around a node $G$ is performed should be by asking an LLM to perform various ‘mutations’ around $G$, and finding the $\alpha$-links among those.
Following this principle, it is tempting to think of the exploration of the space of games similar to ‘random web browsing’ (following the philosophy of the PageRank algorithm [PBRW99]). If in some sense one wants to find a ‘good sample’ of the internet, one can browse randomly and save the visited pages on the fly. Depending on the application we may even move for long enough so our visit becomes independent from the starting point (note also that the longer we ‘walk’ the more biased towards the more connected nodes we get). From this, we could in principle use various heuristics (e.g. the PageRank algorithm) to rank the games that we care about for a target application.
In order to implement the above general idea of ‘exploration’, we still need to clarify what ‘application’ we aim for (at least, implicitly), in particular we try to estimate general capabilities of models. In Section 6.2 below, we propose a principled way to perform this, based on evolutionary ideas.
6.1.2. Toy Model
As a means to gain some intuition on the geometry of the Xent Game space, we can use the following model:
  • Represent by indices $1,2,3,\ldots,n,\ldots$ an (infinite) list of independent ‘skills’ that may be associated with games.
  • Model each game $G$ by an infinite nonzero vector $\left(g_{1},g_{2},\ldots,g_{n},\ldots\right)$ of nonnegative entries that have (uniformly) bounded sum $\|g\|_{1}$, with each entry $g_{i}$ representing naively ‘how much the skill $i$ is involved (and learnable) when playing $g$’.
  • For any pair of games $G,\tilde{G}$, corresponding to $\left(g_{1},g_{2},\ldots,g_{n},\ldots\right)$ and $\left(\tilde{g}_{1},\tilde{g}_{2},\ldots,\tilde{g}_{n},\ldots\right)$, we define $\mathcal{V}_{G}\left(\tilde{G}\right)$ as $\frac{1}{\|\tilde{G}\|_{1}}\sum_{i=1}^{\infty}\min\left(\tilde{g}_{i},g_{i}\right)$, i.e. which fraction of $\tilde{g}$ skills are ‘captured’ by playing $g$ and $\mathcal{V}_{\tilde{G}}\left(G\right)$ as $\frac{1}{\|G\|_{1}}\sum_{i=1}^{\infty}\min\left(g_{i},\tilde{g}_{i}\right)$ the symmetric quantity.
Remark 59.
This simply suggests the (naive, but useful) view that games are associated with skills that are entries in a vector, and that learning these amounts to acquire skills proportionally to the vector of skills.
As we will see in Section 6.3.4, taking random games (by picking a distribution for the skill assignments for games), we can construct a model space to investigate various exploration algorithms numerically.
6.2. Evolution-Based Scope Expansion
6.2.1. Meta-Game: Informal Evolution Dynamics Idea
In a sense, the question of measuring ‘general capabilities’ of models comes down to a motivation: why do we want to evaluate general abilities, in the first place? A plausible answer is because agents will be exposed to unforeseen problems; but then it is important to understand what could possibly generate these unforeseen problems. If we think of challenges that intelligent (animal) agents face in the real world (especially unforeseen ones), they very often come down to the fact that other agents are present (animals are in competition for resources, can be preys or predators, must reproduce, etc.).
These kinds of questions and ideas have been explored in the open-ended evolution context, in particular under the context of Quality-Diversity search (see e.g. [LeSt11a, LeSt11b, Clu19, ECMO23]).
Based on such ideas, we propose to explore the space of Xent Games using a competitive vision which is close in spirit to the Novelty Search with Local Competition (NSLC) algorithms proposed in [LeSt11a, LeSt11b]. We can imagine a world where a number $N$ of agents play a (repeated) meta-game where each move is the generation of Xent Games. The meta-game design elements consist of roughly the following:
  • Agents can choose some games they want to specialize in (i.e. that they decide to learn playing ‘privately’), which they will play with a random sample $k$ of other agents (this would naively correspond to the ‘field’ on which they choose to live) in zero-sum mode.
  • Agents must at the same time play against the $k$ randomly sampled agents (again, in zero-sum mode) the games picked by the latter (on which they will have specialized, by symmetry).
If we think of a competitive landscape, agents will hence be incentivized to pick specialization in games that give them an edge over other agents’ games (because of the zero-sum nature of the games: these are resource allocation games), while at the same time picking games at which other agents’ skills are under-developed. This naturally suggests a competitive dynamics will take place, which should see the games evolve over time.
Remark 60.
This view of an evolving landscape generated by agents by a local game (or meta-game here) is for instance studied in [Lal2015]. The idea of using evolution-related ideas for artificial general intelligence has seen some uptick recently, see in particular [Ada23, Hug24, ZHLLC25]. Note that these ideas are also similar to the ideas of artificial curiosity and self-improvement [Schmi07, Schmi10, Clu19].
Leveraging such a dynamics as a ‘browsing mechanism’ (to use the analogy of Section 6.1.1) to build a game archive (the set of games visited) as an exploration dynamics to measure general abilities is appealing (as far as building a scope is concerned), as it provides a general idea of where ‘unforeseen new games’ could plausibly come emanate from; we could therefore use the list of visited games as a basis for a benchmark.
Remark 61.
While it can definitely be interesting to specify the details of the above meta-game, in the current paper we only use it as a means to inform the selection of relevant games to build appropriate Xent Measures. As such, it is in particular reasonable (assuming universality, see Section 6.2.3 below) to only assume that agents all use the same fine-tuned variants of the same base model $\mathcal{M}_{\mathcal{B}}$.
6.2.2. Key Features and Key Challenges of the Meta-Game
The vision outlined in Section 6.2.1 above promises a number of interesting features:
  • If we start from a set of ‘core abilities’ represented by an initial game $\Sigma_{0}$, we have a good idea of ‘how far we drifted’ from $\Sigma_{0}$ as the process runs, and we can in principle use this to adjust the weights of the new games accordingly.
  • There is fairly clear idea of what we mean by ‘general capabilities’ in this context: the ability to make up for a lack of a specialization, that could realistically be picked and indeed learned by another agent (i.e. for a ‘useful’ reason). In other words, agents that are ‘generally more capable’ will withstand confrontation with other agents on their respective fields, while keeping an edge on their own.
  • It is also consistent with our view on game normalization, where scores represent the ability to make up for a number of attempts: in a sense, we measure the ability to make up for a missed ‘virtual number of attempts’ that could have been emulated by an alternative specialization.
  • If $N$ is not too large, the dynamics will naturally avoid oversampling: agents are incentivized to find ‘original’ directions that are far away from existing games, otherwise they will get ‘eaten’ by competitors; if there is an ‘oversampled region’, it naturally becomes prey for ‘original’ agents.
  • If $N$ is not too small, the agents will at the same time be incentivized to stay ‘generally relevant’, i.e. not to ‘drift towards isolated niches’ (that would deprive them from an ability to compete with the others).
  • This suggests a view about general abilities formed by the aggregate of metastable equilibria: driven by the individual ‘initiatives’ that are selected on the basis being relevant between covering existing abilities and uncovering new directions. Hence the only way from the dynamics to not extend to some more broader (i.e. ‘uncaptured’) region of the game space is if there is no path of ‘meaningful’ games to it.
In spite of the appealing features listed above, a number of complex features are associated with the meta-game described in Section 6.2.1, make it in fact very under-specified and challenging to implement in practice:
  • Many unspecified details about the game are in fact difficult to set: how many players are involved throughout the steps, how are they different from one another, how much can they capitalize upon their previous choices, how are they making their choices, what information do they have about each other, etc?
  • The dynamics suggest an ‘intersubjective’ formulation of general capabilities: general capabilities move in some sense towards a direction resulting from what agents expect general abilities to look like. It can lead to a kind of very unstable set of games that ‘each believes the other believes, but no one truly believes it’.
  • It is not clear how things will behave if the number of players becomes large, and whether determining an optimal strategy is even feasible; if the game is a partial information game, the corresponding Nash equilibria are likely to be impossible to approximate in practice.
  • While this is not necessarily a drawback, the evolution should involve randomness: in natural settings, it is pretty clear that the unpredictable component of evolution is an integral part of it and that no two ‘runs’ of a system will yield the same results.
  • It is not clear exactly how players are supposed to play in practice: should they communicate to cooperate, do they have bounded rationality, should they assume the other players are being played by some fixed models, etc.?
It should be obvious that determining a precise definitive answer to these questions, if one tries to model a real-world-inspired dynamics, is impossible. Ultimately, a way out is possible if the result of the dynamics, as far as measuring general capabilities is concerned, does not depend on these details: this is the view proposed in Section 6.2.3.
6.2.3. Universality: Do Exploration Details Matter?
A way out of the difficulties associated with playing the meta-game discussed above is to ask the question: does the precise meta-game play algorithm really matter? More precisely: if all we care about is selecting lists of Xent Games to benchmark the general abilities of agents, does the specific choice of games picked by the agents in the meta-game actually matter?
Heuristically, there is a very simple reason why specific details should not matter: the whole point of any good measure of general abilities is that it should actually be fairly resilient to changes of the rules of games we use: agents with strong general capabilities will keep their strong performances on games that are known to be connected, while over-specialized agents are unlikely to get systematically lucky over a large, diverse enough family of games. Going further, any strong agent playing the meta-game should be playing in an environment where it can withstand any specific, computable choice of games: any such choice of games is in itself a possible play strategy, and any meta-game agent should choose games so that it can be resilient against such a specific strategy. Similarly, as suggested in Section 4.3.1), we expect the dependence of the measures of general abilities to be fairly independent of the precise setting (the $\mathcal{M},\mathcal{F},\mathcal{T}$ setting) associated with the transfer value functions (as long as they are ‘good’).
It is in fact reasonable to postulate the only relevant features are that the exploration covers a (1) connected and (2) diverse enough web of games, sampled in a (3) fair way. While the connectedness is actually easy to guarantee (by construction), it is important to make sure that the scope does not miss crucial games, i.e. games that are important to learn other games. As we will see in Section 6.3 below, a criterion that essentially implies this is the following:
Assumption 62.
There exists a number $\alpha>0$ and a number $K\geq 1$ such that the space of Xent Games is $\alpha$-connected and such that for any game $G$ the maximal number $n$ of ‘escaping directions’, i.e. of games $G_{1},\ldots,G_{n}$ such that
$\min_{k=1,\ldots,n}\mathcal{V}_{G_{k}}\left(G\right)$ $\geqslant\alpha$
$\max_{j=1,\ldots,k-1}\mathcal{V}_{G_{j}}\left(G_{k}\right)$ $\leqslant 1-\alpha$ $\forall k\geqslant 2$
must satisfy $n\leqslant K$.
Remark 63.
The above assumption suggests it is impossible to find very long sequences of ‘independent’ games that all capture some new meaningful information about any game: in other words, the number of different skills that have a substantial contribution to a game is finite. This natural assumption is for instance easily verified in the toy model introduced in Section 6.1.2 for $\alpha$ small enough and $K=\mathcal{O}\left(\left(1-\alpha\right)/\alpha\right)$.
From the Assumption 62, if there is an ‘undiscovered’ important game, there is at least a chain leading to it, and we will eventually find it, as it is impossible to ‘get lost’ in too many games starting from a game. Note that in practice, it is important to ensure that $K$ is reasonably small. From the Section 6.3, we propose a simple greedy algorithm aimed at extracting a sequence of games which arguably captures the same abilities as those discovered by a dynamics like the one suggested in Section 6.2.1.
Remark 64.
A very short summary of the above discussion is simply: any computable algorithm is a strategy, so any generally capable agent must be ready to face it, and conversely, any strategy must uncover a family of connected games, so any fine-grained enough will uncover them.
6.3. Scope Growth Algorithm
6.3.1. Simple-Minded Point of View
The success of the ideas of the Novelty Search with Local Competition algorithms together with universality considerations postulated above (Section 6.2.3) suggest that naive algorithms for evolutionary processes like the ones discussed in Section 6.2.1 are a promising way to search for an extended Xent Game scope.
6.3.2. Algorithm Specification
From the considerations made in Section 6.2 above, we propose the greedy scope growth algorithm, which builds an archive of games (in the terminology of the Quality-Diversity algorithms) as a means to build Xent Measures.
  • Parameters: $\alpha_{1}>0,\alpha_{2}>0$, a meta-sampling process $\Lambda$ (see Section 6.3.3).
  • Start with an initial scope $\Sigma_{0}$.
  • For $k=0,1,2,\ldots,N$
    • Set $\Omega_{k}=\emptyset$.
    • For each $G$ in a randomized ordering of $\Sigma_{k}\setminus\Sigma_{k-1}$:
        Repeat until stop:
          Sample from the meta-sampling process $\Lambda$ a list of games $\mathcal{G}_{j}$ related to $G$.
          Find the $G_{*}\in\mathcal{G}_{j}$ such that $G_{*}\rightmoon_{\alpha_{1}}G$ that achieves the smallest value of $\gamma:=\max_{\tilde{G}\in\Sigma_{k}}\mathcal{V}_{\tilde{G}}\left(G_{*}\right)$ possible.
          If $\gamma\leq 1-\alpha_{2}$, add $G_{*}$ to $\Omega_{k}$.
          Else: stop.
    • Set $\Sigma_{k+1}=\Sigma_{k}\cup\Omega_{k}$ if $\Omega_{k}\neq\emptyset$.
    • Output the Xent Measure $\Sigma_{k+1}$.
    • Stop if $\Omega_{k}=\emptyset$.
Algorithm 1 Scope Growth Algorithm
Remark 65.
By the Assumption 62, each ‘repeat’ round is guaranteed to stop after $K$ steps. In practice, the fact that the list of games $\mathcal{G}_{j}$ must be sampled from an LLM will make this stop much faster.
From our picture of the Xent Game space, we obtain the following:
Claim 66.
From Assumptions 55 and 62, for small enough $\alpha>0$, the Scope Growth Algorithm (1) with $\alpha_{1},\alpha_{2}\leq\alpha$ covers the space of playable Xent Games as $N\to\infty$.
Remark 67.
While we do not have explicit fairness guarantees, Algorithm 1 tends to avoid the oversampling of any region (as any new game must avoid being $\left(1-\alpha\right)$-covered by the previous games), thus limiting the oversampling risk.
6.3.3. Meta-Sampling Process
It must be noted that while the scope growth algorithm presented above is very simple in structure, much of the implementation is left in the choice of the meta-sampling process $\Lambda$.
The objective of $\Lambda$ is quite clear: to provide the largest possible sample of games that allows the loop of the scope growth algorithm to run for as long as possible, in a way that maximally exhausts games. Given the structure of the Xent Games, the most useful practical implementations of such an algorithm involve sampling random codes in the XGL language from a dedicated model, trained to maximize the sampling quality. As discussed in Section 2.4.1, the Xent Game structure and the XGL is designed from the beginning with this goal in mind.
6.3.4. Numerical Simulation: Exploration in Toy Model
Taking the toy model defined in Section 6.1.2, we can illustrate the scope growth algorithm in a naive 2D setting (representing games as uniform random positive points under a curve $\sqrt{x}+\sqrt{y}=1$), by taking the meta-sampling $\Lambda$ to give the list of points within a small neighborhood of radius $\rho$ of the game $G$. As expected, the algorithm progressively crawls towards the Pareto efficiency region, and ultimately covers it completely.
Figure 6.1. Scope growth algorithm exploration steps $2D$ with $\alpha_{1}=0.9$, $\alpha_{2}=0.05$, $\rho=0.1$.
7. Summary, Discussion and Perspectives
7.1. Summary
Informally speaking, this paper follows a path from implicit knowledge to general capability measures:
  • We start from the question of implicit knowledge, and formulate a number of concrete tasks associated with it.
  • We propose to approach these tasks using a common game-theoretic framework, and ask the question of what is a good space of games to work with.
  • We introduce Cross-Entropy Games (Xent Games) as an answer to the previous item, characterize this space theoretically and provide examples of tasks covered by it.
  • We investigate the question of what it means for an LLM to play Xent Games, propose evaluation procedures and introduce notions of playability and of transfer value.
  • We propose to evaluate LLMs by making them play Xent Games, propose a score normalization scheme for playable games, and, given a scope (a set of playable games), propose to construct Xent Measures by considering optimal covers in terms of transfer value.
  • We propose to address the challenge of unbounded scope associated with measuring general capabilities (as opposed to limited to a finite scope) by relying on evolutionary dynamics principles, leading us to considerations close in philosophy to Quality-Diversity ideas.
  • Motivated by universality considerations, we propose a simple greedy scope growth meta-algorithm as a general capability measure.
7.2. Discussion
A number of important points in the approach above deserve some discussion.
7.2.1. Role of Judge Model
Throughout the paper, the Judge Model $\mathcal{J}$ is taken as a given. A central point is that for a sufficiently strong judge model, the space of implicit knowledge questions becomes large enough so that games can be played ‘ad infinitum’ (similarly to the fact that once one knows the rules of chess, one can keep improving at the game for a very long time). Still, improving the abilities of $\mathcal{J}$ or (perhaps counter-intuitively) allowing $\mathcal{J}$ to be ‘controllably weak’ (e.g. with certain prefix tokens) can lead to more sophisticated games; ultimately, the role of $\mathcal{J}$ is to provide an environment where abilities ought to be developed; as such, having high-quality data from the ‘external world’ contained in the training of the $\mathcal{J}$ model is fundamental, as the $\mathcal{J}$ model is the only connection between the external world and the model playing games.
7.2.2. Units of Measure
An important theme of Sections 4, 5, and 6 is to provide a clean interpretation of the quantities computed, in particular good units of measure. While the raw Xent Games values are measured in bits (which carries some useful problem-specific information), the measures shift towards relative units of work for the score normalizations with respect to a base model (comparing number of attempts) and the transfer value (such values are normalized by comparing token processing counts).
Ultimately this leads to a simple perspective on general capabilities which is in terms of units of work spared compared to specific replaying and training, assuming a common ‘reality’ defined by a model $\mathcal{J}$. As such, this provides a view of general capabilities that is continuous and does not contain any explicit thresholds for capabilities.
7.2.3. Connections
Ultimately, our construction relies upon a combination of ideas which appear in information theory, compositional game theory, transfer learning, search, benchmarking, and open-ended evolution. While some combinations of the above ideas are definitely well-studied, it is interesting to note that to the best of our knowledge, the connection between compositional game theory and transfer learning has not been studied at all; it is also interesting that the connection between open-ended evolution and general intelligence has seen a recent surge in interest [Ada23, XDCGZ24, ZHLLC25], but that no connection with the question of measuring intelligence has been provided yet.
7.3. Perspectives
Building the foundations of the universality ideas (Section 6.2.3) appears to be one of the most interesting theoretical challenges associated with our results: while usually the evolutionary ideas are often proposed as a very imperfect means towards an end, it is possible that as far as building general capability measures via meta-games, the imperfection disappears for game-theoretic reasons.
7.3.1. Implementation
The most enticing practical challenge associated with the vision proposed in this paper is its implementation at scale: the determination and study of families of Xent Measures generated according to the ideas proposed in Section 6, including in particular efficient means to perform the game meta-sampling. In parallel, an interesting question is to investigate the potential of hand-crafted Xent Measures (i.e. for a finite set of hand-crafted games) as restricted benchmarks: in spite of their inherently narrow scopes, these already satisfy a number of interesting properties (in particular, not relying on a private dataset, and being generally hard to overfit on) that make them promising contenders to measure the abilities of models.
7.3.2. Synthetic Data
An important question in the field of Language Models is that of synthetic data (see e.g. [WWZ24, LWLZD24]) for LLM training. An interesting question is to determine how much value can be learned by pre-training LLMs on synthetic data coming from gameplay on Xent Games. The long-context window and other challenging gameplay aspects could in principle lead pre-trained models to learn deeper and more nuanced representations of information, enabling them to capture subtle notions associated with the implicit knowledge tasks.
7.3.3. Curriculum Learning
An interesting direction is the construction of curriculum-learning algorithms: the approach we have proposed so far presupposes the presence of agents whose abilities are to be measured, but leaves aside the question of how such abilities ought to be discovered. A promising path is to leverage evolutionary ideas such as the ones outlined in the present paper to build such curricula, i.e. to identify the most relevant games to learn to play in order to develop general abilities.
7.3.4. Putting Things Together
All in all, in the setting of Xent Games, we have LLMs play several distinct roles. Besides the base models, used for normalization, we have LLMs:
  • As judge models
  • As models for the NPC players in the games
  • As models to generate the game maps
  • As meta-sampling models to generate games
In principle, these models can all be the same, and they can evolve over time. If one leverages Xent Games as means to train new models (as suggested in Sections 7.3.2 and 7.3.3 above), then this can lead to a self-improvement loop of LLMs, which is an interesting alley to pursue: this is one way in which open-endedness ideas could (in principle) lead to an augmentation of capabilities for LLMs in new dimensions.